### v3

# Meta-Reasoner: Dynamic Guidance for Optimized Inference-time Reasoning in Large Language Models

Abstract
Large Language Models (LLMs) increasingly rely on prolonged reasoning chains to solve complex tasks. However, this trial-and-error approach often leads to high computational overhead and error propagation, where early mistakes can derail subsequent steps. To address these issues, we introduce Meta-Reasoner, a framework that dynamically optimizes inference-time reasoning by enabling LLMs to “think about how to think.” Drawing inspiration from human meta-cognition and dual-process theory, Meta-Reasoner operates as a strategic advisor, decoupling high-level guidance from step-by-step generation. It employs contextual multi-armed bandits to iteratively evaluate reasoning progress and select optimal strategies (e.g., backtrack, clarify ambiguity, restart from scratch, or propose alternative approaches), and reallocates computational resources toward the most promising paths. Our evaluations on mathematical reasoning and puzzles highlight the potential of dynamic reasoning chains to overcome inherent challenges in the LLM reasoning process and also show promise in broader applications, offering a scalable and adaptable solution for reasoning-intensive tasks.
Meta-Reasoner: Dynamic Guidance for Optimized Inference-time
Reasoning in Large Language Models
Yuan Sui1, Yufei He1, Tri Cao1, Simeng Han2, Yulin Chen1, Bryan Hooi1 1 National University of Singapore 2 Yale University yuan.sui@u.nus.edu, bhooi@comp.nus.edu.sg
1 Introduction
Recent advances in o1-like and r1-like [^1] O1 (from OpenAI) and R1 (from Deepseek) are two advanced LLM models designed for reasoning. reasoning have enabled large language models (LLMs) to achieve remarkable performance on complex tasks such as mathematics Patel et al. (2024); Lightman et al. (2023), science Rein et al. (2023), and logical puzzles Lei et al. (2024); Yao et al. (2023). By simulating multi-step, human-like deliberation Yao et al. (2024), these methods allow LLMs to decompose problems into smaller subproblems, test hypotheses, reflect on intermediate results, and iteratively refine their solutions. This extended reasoning process enables systematic exploration of ideas, verification of partial conclusions, and progressive improvement before producing a final answer. Such capabilities are particularly valuable in domains demanding rigorous logical reasoning Chenghao Yang (2024).
Despite these advances, o1/r1-like reasoning remains fundamentally challenged by its trial-and-error nature: models generate numerous candidate reasoning paths, discard flawed ones, and gradually converge on solutions. While this flexibility facilitates exploration of diverse strategies, it often incurs substantial computational overhead Snell et al. (2024); Manvi et al. (2024) and is vulnerable to error propagation, where early mistakes accumulate and compromise subsequent steps Lei et al. (2024); Yao et al. (2023); Gandhi et al. (2024). Some iterative methods incorporate partial revision or backtracking Gandhi et al. (2024); Li et al. (2025a), but these approaches tend to be ad-hoc and limited to correcting errors within a narrow reasoning window. Crucially, they lack a systematic mechanism to assess whether an entire reasoning trajectory remains promising or should be abandoned.
As a result, LLMs risk becoming “stuck” on unproductive reasoning paths, wasting valuable computational resources without recognizing when a strategic pivot is necessary. A critical challenge, therefore, is to enable LLMs to manage their reasoning budget more effectively—prioritizing promising directions while adapting or discarding ineffective strategies during inference time.
To address this challenge, we propose Meta-Reasoner, a specialized meta-reasoning module that operates alongside the LLM to enhance its reasoning capabilities. Acting as a high-level advisor, the meta-reasoner dynamically evaluates the reasoning process and provides strategic guidance or redirection when progress stalls. Unlike the LLM, which focuses on detailed stepwise generation, the meta-reasoner maintains a global perspective, assessing overall progress and strategy from a high level. Meta-Reasoner operates in iterative rounds: First, the LLM generates partial chain-of-thought reasoning chains and a concise “progress report” summarizing its current state. Meta-Reasoner then reviews this report and offers high-level feedback - such as restarting reasoning with a different approach, refining existing ideas, or focusing on specific subproblems. This setup allows Meta-Reasoner to concentrate on overall strategy rather than getting involved in the granular details of the LLM’s reasoning. Overall, Meta-Reasoner helps prevent the LLM from getting stuck or spending resources on unproductive lines of inquiry during the inference time.
Overall, the main contributions are as follows:
  • We propose a meta-reasoning module that provides high-level oversight and guidance to the LLM’s reasoning, helping to prevent it from getting stuck on unproductive solution paths during inference.
  • We design a lightweight progress reporting mechanism, enabling efficient communication of the LLM’s reasoning state to the meta-reasoner with minimal overhead.
  • We empirically show that our Meta-Reasoner improves accuracy and efficiency on challenging math and scientific reasoning benchmarks (e.g., Game of 24, TheoremQA, and SciBench) compared to existing strong baselines. Our results highlight the framework’s potential as a new scalable solution to inference-time reasoning.
2 Related Works
In this section, we first review the challenges of complex reasoning in LLMs and the limitations of chain-of-thought (CoT) methods. We then discuss recent advances in backtracking and self-verification techniques that aim to mitigate these issues. Finally, we draw inspiration from cognitive science, framing our approach within the meta-cognition and dual-process theory to motivate the design of our Meta-Reasoner framework.
Complex Reasoning in LLMs
The introduction of CoT reasoning has transformed how LLMs tackle complex problems by decomposing tasks into intermediate steps Lee et al. (2025). Recent models such as OpenAI’s o1 and o3, and Deepseek’s r1, have achieved state-of-the-art performance across diverse domains by leveraging CoT-like reasoning Manvi et al. (2024); Li et al. (2025a); Kudo et al. (2024); Sui et al. (2024b). However, the inherent sequential dependency in CoT limits robustness: errors in early steps can propagate and degrade overall performance Snell et al. (2024). Moreover, when confronted with complex reasoning tasks, CoT methods may become trapped in infinite reasoning loops Lee et al. (2025); Sui et al. (2024a, c). To address these challenges, we propose Meta-Reasoner, a meta-reasoner that monitors and adapts the reasoning strategy based on the progress of CoT. By integrating mechanisms such as backtracking and self-verification, Meta-Reasoner provides a holistic perspective that evaluates the overall reasoning trajectory to prevent stagnation and reduce wasted computation on unproductive lines of thought.
Backtracking and Self-Verification
To mitigate the limitations of CoT-like reasoning, recent methods have explored backtracking and self-verification techniques Yao et al. (2023); Besta et al. (2023); Gandhi et al. (2024). For example, Weng et al. (2023) demonstrate that incorporating a self-verification stage—where the model re-examines its conclusions using the generated chain of thought—significantly improves performance by detecting errors early. Similarly, Ling et al. (2023) propose generating multiple candidate reasoning chains alongside a verifier mechanism that identifies and backtracks on erroneous steps. These approaches extend beyond post-hoc validation by enabling dynamic strategy adjustments during inference Lightman et al. (2023), thereby limiting error propagation in lengthy reasoning chains and mitigating infinite reasoning loops. Building on these efforts, our Meta-Reasoner framework employs instructions to (1) restart from scratch with alternative strategies, (2) backtrack to the error point, and (3) continue with targeted suggestions. Further details on this strategy are provided in §4.3.
Meta-Cognition & Dual-Process Systems
From a cognitive science perspective, meta-cognition involves higher-order processes that allow individuals to monitor, evaluate, and adjust their cognitive strategies Gao et al. (2024); Yoran et al. (2024). This reflective thinking—often characterized as System 2 in dual-process theories Havrilla et al. (2024)—is vital for tasks requiring careful deliberation and error correction Didolkar et al. (2024). Drawing on these insights, our Meta-Reasoner framework can be viewed as analogous to dual-process systems: the LLM generates CoT steps akin to System 1, while the Meta-Reasoner provides high-level strategic oversight, analogous to System 2, guiding or redirecting reasoning as needed. This separation of responsibilities balances efficiency with robust problem-solving, allowing the LLM to handle routine inferences and the Meta-Reasoner to intervene for strategic adjustments.
3 Preliminary
In complex reasoning tasks, a key challenge is deciding which strategy to use from multiple valid options. This problem can be naturally framed as a contextual multi-armed bandit (MAB) problem, a well-studied framework for making decisions that balance trying new options (exploration) and using what is already known to work (exploitation).
Imagine an agent faced with several strategies (called arms), and at each step, it observes some information about the current situation (called the context). Based on this context, the agent picks one strategy to apply and then receives feedback (a reward) indicating how well that strategy performed. The goal is to choose strategies over time to maximize the total reward.
Formally, at each time step $t$, the agent observes a context vector $x_{t}$ describing the current state and selects an arm $s_{t}$ from a set of possible strategies $\mathcal{S}$. After choosing $s_{t}$, it receives a reward $r(s_{t},x_{t})$ that depends on both the chosen strategy and the context. The agent aims to maximize the cumulative reward over $T$ steps:
$R(T)=\sum_{t=1}^{T}r(s_{t},x_{t}).$ (1)
A popular algorithm for this problem is LinUCB Li et al. (2012), which assumes the expected reward is approximately a linear function of the context. Specifically, for each arm $s$, there is an unknown parameter vector $\theta_{s}$ such that:
$\mathbb{E}[r(s,x_{t})]\approx x_{t}^{\top}\theta_{s}.$ (2)
LinUCB maintains an estimate $\hat{\theta}_{s}$ of this parameter and a measure of uncertainty about the estimate. At each step, it selects the arm that maximizes a combination of the estimated reward and an uncertainty bonus:
$s_{t}=\arg\max_{s\in\mathcal{S}}\left[x_{t}^{\top}\hat{\theta}_{s}+c\sqrt{x_{t}^{\top}A_{s}^{-1}x_{t}}\right],$ (3)
where $A_{s}$ is a matrix capturing past observations for arm $s$, and $c$ controls how much the agent favors exploring uncertain options. The second term encourages trying arms that might perform well but have been less explored.
By using context to guide its choices, LinUCB adapts its strategy selection to the current situation, which aligns with our goal in Meta-Reasoner: to dynamically choose the most effective reasoning strategy during inference time.
4 Methods
Motivated by the intuition that LLMs should concentrate their computational efforts on more promising reasoning paths during inference time, we explore two key research questions in this paper: (1) How can language models dynamically allocate resources during inference to optimize reasoning and planning?; (2) What architectural design enables an effective separation between the reasoning process within the LLM and the meta-level guidance that oversees it?
To address these questions, we propose a novel framework, Meta-Reasoner, which endows LLMs with the ability to “think about how to think”. Our framework supervises the reasoning process and dynamically guides the model to focus on more promising reasoning trajectories during inference time. Furthermore, Meta-Reasoner mitigates the limitations of conventional sequential reasoning, which may get stuck in suboptimal paths. We propose a “high-order" reasoning mechanism to balance exploration and exploitation using a Multi-Armed Bandit (MAB) algorithm.
Figure 1: An illustration of the Meta-Reasoner workflow. In each round, the LLM produces a new reasoning step to extend its CoT reasoning. The CoT is then summarized into a progress report, which provides context for the meta-reasoner. Then meta-reasoner uses a contextual multi-armed bandit (either using a fixed contextual bandit or dynamic contextual bandit) to choose a guidance strategy. The selected strategy then guides the next reasoning step generation, to enable strategic redirection, error correction, and resource optimization. A reward is then computed from the progress report and used to update the bandit algorithm. The process repeats until the task is complete or the maximum number of rounds is reached.
The meta-reasoning framework operates iteratively as illustrated in Figure 1. At each round $t$, the reasoning process comprises three steps: (1) CoT generation by the LLM, (2) Progress Reporting to summarize the reasoning progress so far (i.e., this is partly for efficiency, and partly to help the meta-reasoner focus on its main goal of “advising” rather than being distracted by the details in the CoT), and (3) Strategy Generation by the meta-reasoner to optimize subsequent steps. The selection of the strategy is almost exactly corresponds to the well-studied problem of contextual multi-armed bandits as illustrated in §3. Each strategy can be seen as an arm for the bandit, and the reward of each strategy can be evaluated by the progress of LLM reasoning after applying the strategy. We analogy the process of executing and evaluating each strategy as the act of “pulling” each arm. The overall goal of our meta-reasoner is to find the best arm (i.e., strategy with highest cumulative rewards) with as few pulls as possible. The complete algorithm of Meta-Reasoner is appended in Algorithm 1.
4.1 Chain-of-Thought (CoT) Generation
In the first step, the LLM generates a reasoning step to extend its CoT reasoning based on the user query. Starting from its reasoning history $C_{t-1}$ and the guidance $G_{t-1}$ provided by the meta-reasoner in the previous round, the LLM $M$ produces a new reasoning step $s_{t}$. This step is appended to the current CoT, forming $C_{t}=C_{t-1}\cup\{s_{t}\}$. By maintaining the full reasoning trajectory at each round, the model establishes a coherent foundation for evaluation and further refinement. This approach resembles the long-term reasoning demonstrated in models like o1, which generate extended CoTs. However, this reasoning process often resembles “trial-and-error”, potentially incurring unnecessary inference costs on unproductive paths. Moreover, due to its sequential nature, the process is prone to becoming trapped in suboptimal solutions.
4.2 Progress Reporting
After updating the CoT, we summarize the reasoning history $C_{t}$ into a concise progress report $P_{t}$. This summary captures key aspects of the reasoning trajectory, including progress toward the task goal, reasoning consistency, and significant updates thus far. The summarization function $f$ abstracts the detailed CoT into a simpler, more focused representation. This step is designed to be both computationally efficient and informative, enabling the meta-reasoner to evaluate high-level progress without being overwhelmed by the granular details of every reasoning step. While this may be viewed as an engineering heuristic, we observe that including essential information in the prompt unlocks the LLM’s capacity for “higher-order” thinking. Consequently, the LLM tends to generate more insightful and critical strategies, which are particularly valuable for complex reasoning tasks.
Diagnosis Strategy
Progress is insufficient or the current strategy seems ineffective. Restart from scratch and propose alternative strategies.
[…]
There are mistakes in intermediate steps. Backtrack to the point where the error occurred.
[…]
The current approach is working well. Continue and provide specific suggestions for the next steps.
Ambiguous or conflicting intermediate results are observed. Pause to clarify and disambiguate the current reasoning, then reconcile the discrepancies.
[…]
The reasoning process appears overly complex or convoluted. Simplify by decomposing the task into smaller, manageable sub-tasks.
[…]
Evidence of error propagation or low confidence in certain sub-components. Perform targeted verification on critical steps and focus on areas with low confidence.
[…]
Repetitive or circular reasoning patterns are detected. Reset to a previously successful checkpoint and explore alternative solution paths.
Table 1: Demonstration: Contextual bandit pair (i.e., diagnosis of the progress report (context) and the corresponding strategy (bandit)) for guiding the LLM’s reasoning process. Marked rows are some of the unique strategies generated by Dynamic Contextual Bandits.
4.3 Meta-reasoner Strategy Generation
In the next step, the meta-reasoner evaluates the progress report $P_{t}$ and selects an appropriate strategy $G_{t}$ for LLM reasoning (the complete procedure is detailed in Algorithm 1). We formulate the generation of strategy as a contextual MAB problem and consider two settings: (1) a fixed-strategy formulation, where the meta-reasoner selects from a predefined set of strategies using a contextual bandit algorithm; and (2) an advanced setting, in which the meta-reasoner itself is an LLM-based agent capable of introducing or refining strategies dynamically. In both cases, the meta-reasoner employs the partial-feedback principle of MABs to adaptively select strategies based on a reward function that evaluates the quality of reasoning progress after applying the chosen strategy. We demonstrate the contextual bandit pair (i.e., diagnosis of the progress report (i.e., context) and the corresponding strategy (i.e., bandit) in Table 1.
Progress Evaluation.
A central objective of our evaluation mechanism is to quantify how effectively the model’s current reasoning advances toward the task goal (e.g., solving a complex problem), while also monitoring computational cost to promote efficiency. Concretely, we implement a reward function that tracks both solution progress (e.g., partial correctness, adherence to constraints) and resource usage (e.g., number of reasoning steps). This evaluator can be any suitable mechanism, including LLM-based verification or external scoring scripts. In our implementation, we leverage an LLM as the evaluator (with prompts referenced in Figures 47) to assess reasoning progress. The evaluator outputs a cumulative score, which is then used to update the MAB algorithm.
Fixed Contextual Bandit.
In the basic version of our framework, the meta-reasoner is modeled as a single contextual bandit that selects from a fixed, finite set of $K$ strategies. These strategies include instructions such as “continue and provide specific suggestions”, “restart from scratch”, “backtrack to the point where the error occurred”, or “propose alternative methods or perspectives to consider”, as detailed in Table 1 (rows without marked). At each round, the LLM produces a progress report summarizing its partial reasoning, the meta-reasoner transforms this progress report into a feature vector $x_{t}$ using a language model and applies a contextual bandit algorithm (e.g., LinUCB Li et al. (2012)) to select the best next strategy $a_{t}$. The LLM then executes that strategy and we collect the reward $r_{t}$ for $a_{t}$ based on the reward function. Through iterative MAB algorithm updating, the MAB algorithm learns to select appropriate strategies conditioned on the recent progress report.
Dynamic Contextual Bandit.
The fixed-arm formulation assumes a static set of strategies. In practice, the meta-reasoner may itself be an LLM capable of inventing new strategies over time. To accommodate dynamic strategies, we allow the meta-reasoner to propose or refine new strategies at round $t$, which generates an expanding collection of strategies, $G_{1}\subseteq\cdots\subseteq G_{t}$. Each newly proposed strategy becomes an arm in the contextual bandit. To encourage at least some exploration on this new arm, we initialize each arm with a blank or weak prior in the bandit’s parameters. We further analyze the stability of these new generated dynamic contextual bandits in Appendix §C.
By explicitly separating low-level content generation (handled by the LLM) from high-level strategy decisions (governed by the meta-reasoner’s bandit), the system can effectively avoid getting stuck or wasting excessive resources on poor solution paths. In domains where a predefined set of strategies is sufficient, the fixed-arm formulation can simplify the method deployment. While in more open-ended domains where novel tactics may emerge, dynamic-arm extensions give meta-reasoner more flexibility to evolve.
5 Experiments
In this section, we first introduce the experiment settings including datasets, backbone models and training details of MAB. We then present the main results of Meta-Reasoner with analysis regarding efficiency, rewards accumulation, and qualitative assessment of meta-reasoner output. We present the detailed baselines in Appendix A.
5.1 Experiments Setup
Datasets.
We test our method on several challenging datasets that demand complex reasoning and often involve lengthy thinking processes for the correct solutions. These includes (1) 24-point game Yao et al. (2023); (2) college-level scientific problem from SciBench Wang et al. (2024) and (3) math questions based on theorems from TheoremQA Chen et al. (2023). For SciBench, we focus only on the math-related subsets (i.e., diff, stat, and calc). Detailed explanations for each subset can be found in Wang et al. (2024). For TheormQA, we only consider the math subset that involves logical reasoning.
Training Details.
We collect the training data for each task: (1) for the 24-point game, we random sample 50 queries from 4nums.com specifically excluding problems in ranks 901-1000 which were reserved for testing; (2) for TheoremQA, we randomly sample 30 mathematical reasoning queries from the dataset; (3) for SciBench, we randomly sample 30 queries from differential subsets including diff, stat, and calc from the entire dataset. These samples were used to iteratively update the LinUCB parameters for both fixed ($K$=3 or $K$=5 strategies) and dynamic strategy settings.
We configure the training process using deterministic generation ($n$=1, Top_k=1, temperature=0) with specific max_token limits for CoT generation (512), meta-reasoner feedback (256), progress reports (512), and reward model outputs (4). The LinUCB exploration parameter was set to $c$=0.2. Experiments were ran for a maximum of $T$=30 iterations for the 24-point game, scibench, and $T$=100 for TheoremQA, with MAB parameters updated after each iteration based on a reward function that weighted objective completion (40%), progress quality (30%), efficiency (15%), and strategy alignment (15%). The reward function prompt is detailed in Figure 4-7.
Method Diff(%) Stat(%) Calc(%)
Phi-4 + CoT 17.42 28.42 32.93
Llama-3.1-instruct + CoT 33.14 49.72 54.18
Gemini-Exp-1206 + CoT 36.32 56.73 59.24
Gemini-Exp-1206 + SC-CoT 38.73 59.12 64.11
GPT-4o-mini + CoT 33.12 55.71 58.10
GPT-4o-mini + SC-CoT 37.33 56.67 63.81
GPT-4o-mini + MCR 40.12 58.21 67.42
[…]
GPT-4o-mini + MACM Lei et al. (2024) 54.78 67.13 65.77
GPT-4o + MACM Lei et al. (2024) 61.42 78.32 76.72
GPT-4o-mini + Meta-Reasoner (our work) 60.32 73.64 80.23
GPT-4o + Meta-Reasoner (our work) 67.14 83.29 84.17
Table 2: Accuracy (%) comparison of different methods on the math-related subset of the SciBench dataset. Each column refers to the problem subset defined in Wang et al. (2024).
Method Accuracy (%)
GPT-4o-mini + CoT Yao et al. (2023) 4
GPT-4o-mini + SC-CoT Yao et al. (2023) 9
GPT-4o-mini + IO (best of 100) Yao et al. (2023) 33
GPT-4o-mini + CoT (best of 100) Yao et al. (2023) 49
Gemini-Exp-1206 + IO (best of 100) Yao et al. (2023) 38
Gemini-Exp-1206 + CoT (best of 100) Yao et al. (2023) 60
[…]
GPT-4o-mini + ToT ($b=1$Yao et al. (2023) 32
GPT-4o-mini+ ToT ($b=5$Yao et al. (2023) 65
GPT-4o-mini + Reflexion Shinn et al. (2024) 53
GPT-4o-mini + MACM Lei et al. (2024) 80
GPT-4o-mini + Meta-Reasoner (our work) 89
GPT-4o + Meta-Reasoner (our work) 92
Gemini-Exp-1206 + Meta-Reasoner (our work) 94
[…]
o1-mini + IO 89
o1-preview + IO 93
Table 3: Accuracy(%) comparison of different prompting methods on 24-points game. $b$: Search breadth.
Backbone Models.
We consider both LLMs and the recent Large Reasoning Models (LRMs) for our experiments. For the LLMs, we consider the closed-source models like gpt-4o, gpt-4o-mini (between Nov 2025 to Jan 2025) from OpenAI, and open-sourced models like meta-llama-3.1-8B-instruct from Meta, phi-4 from Microsoft and gemini-experimental-1206 from Google. For the LRMs, we consider the closed-source models like o1, o1-mini (In case we cannot break down the generation of o1 models through APIs, we cannot properly inject our meta-reasoner with o1-series models; we only provide the IO results for references). For the feature extraction mentioned in §4.3, we use text-embedding-3-small from OpenAI as the embedding model.
To ensure the reproducibility of the experiments, we set $\mathrm{temperature}=0.7$ and $\mathrm{top\_p}=1.0$ for all models. We use the API service from OpenAI [^2] https://openai.com/ and OpenRouter [^3] https://openrouter.ai/ for our experiments which host detailed snapshots of the utilized model versions.
5.2 Main Results
We compare the accuracy of different prompting methods across different backbone models on SciBench (as shown in Table 2), 24-points game (as shown in Table 3) and TheoremQA (as shown in Table 4). We found that basic prompting strategies, such as CoT and SC-CoT, show limited effectiveness, achieving only 4% and 9% accuracy on 24-point games, respectively. Incorporating IO strategy with “Best of 100” samples improves accuracy to 33%, but it remains far behind advanced methods. Strategies like ToT illustrate the importance of exploring broader reasoning paths, with accuracy increasing from 45% to 74% as the search breadth expands from 1 to 5. Advanced iterative methods, such as Reflexion (53%) and MACM (80%), further demonstrate the value of refined reasoning frameworks. Our proposed Meta-Reasoner outperforms these approaches, achieving 89% accuracy with GPT-4o-mini and 92% with GPT-4o, showcasing its ability to dynamically guide reasoning, correct errors, and focus resources effectively. Compared to specialized models like o1-mini, our method equipped with much cheaper and generalized models like GPT-4o-mini delivers comparable performance, demonstrating its adaptability and scalability. Overall, the Meta-Reasoner framework provides a compatible approach to improving reasoning-intensive tasks, combining high accuracy with dynamic and efficient problem-solving strategies. The results on SciBench and TheoremQA also demonstrate similar findings and show that Meta-Reasoner generally achieves better performance compared to the baselines and the results are consistent across different models.
5.3 Ablation Study
Method Accuracy (%)
GPT-4o-mini + CoT 39.46
Gemini-Exp-1206 + CoT 43.12
[…]
GPT-4o-mini + Reflexion Shinn et al. (2024) 74.32
GPT-4 Turbo + MACM Lei et al. (2024) 79.41
GPT-4o-mini + Meta-Reasoner (our work) 84.13
Gemini-Exp-1206 + Meta-Reasoner (our work) 86.32
Table 4: Accuracy(%) comparison of different prompting methods on TheoremQA Chen et al. (2023).
Model Variant Game-of-24(%) TheoremQA(%)
GPT-4o-mini Full Method 89 84.13
w/o Progress Report 85 79.42
w/o MAB (direct arm selection) 82 80.74
w/o MAB (CoT) 4 39.46
Gemini-Exp-1206 Full Method 94 86.32
w/o Progress Report 91 81.78
w/o MAB (direct arm selection) 87 82.14
w/o MAB (CoT) 11 43.12
Table 5: Ablation study of Meta-Reasoner. Direct arm selection refers to prompting LLM to directly select a strategy based on recent progress report.
Bandit Type Game-of-24(%) #US TheoremQA(%) #US
Fixed (K=3) 65 3 72.34 3
Fixed (K=5) 72 5 79.17 5
Dynamic 89 14 84.13 21
Table 6: Fixed vs. Dynamic Bandit Variants over GPT-4o-mini. #US: Number of Unique Strategies.
Figure 2: We normalize the inference costs (based on token usage) in a range from 0 to 1 for easier comparison. We use gpt-4o-mini as the backend model for all the methods in this figure. For each method, key hyper-parameters (e.g., ${N}$ in Best of N, or tree size in ToT) are tuned to yield a baseline (lower point) and an extended (upper point) configuration, with dashed lines connecting these bounds.
In this section, we conduct an ablation study to analyze each component contribution of Meta-Reasoner. In specific, we consider the following setup: (1) w/o progress report: we replace the progress reporting process with directly considering the entire CoT history without summarization; (2) w/o MAB: instead of using MAB to select the proper strategy, we directly leverage an LLM to the decision making to provide the proper strategy for LRM reasoning. In Table 5, we show that when removing progress reporting (“w/o Progress Report”), the overall performance moderately degrades and we hypothesize it is due to the concise intermediate summarizations can help the Meta-reasoner only consider the high-level strategy instead of being confused with too much details of the reasoning process. We also find that removing the MAB brings a more pronounced effect, especially when strategy selection falls back to a direct chain-of-thought approach (“w/o MAB (CoT)”). It verifies the effect of our meta-reasoner module to help the model stay on track for getting an optimal solution. In Table 6, we compare fixed and dynamic bandit variants on the game of 24 and theoremQA. We find that using a fixed set of strategies (e.g., $K=3$ and $K=5$) yields lower performance compared to the dynamic approach which adaptively explores more strategies (shown by larger unique strategies). The results highlight the benefit of flexibly allocating diverse reasoning strategies using LLM in-context learning capabilities.
5.4 Analysis
Figure 3: Cumulative reward of different settings across iteration. We compare our method using LinUCB with baseline (direct arm selection), and random search methods across two tasks—Game of 24 (top row) and TheoremQA (bottom row) using GPT-40-mini (left) and Gemini-Exp-1206 (right).
Inference Efficiency.
A common way to measure computational efficiency is wall-clock time, as it directly reflects real-world latency during inference. However, in our API-based setup, wall-clock time is unreliable due to external factors such as network latency and server load, which introduce noise unrelated to the model’s intrinsic computational cost. Therefore, we use token usage as a more stable and meaningful proxy, since it directly correlates with the amount of computation performed and is widely adopted in recent works Liu et al. (2025); Li et al. (2025b). The detailed instruction can be found in Appendix B.1.
Figure 2 presents normalized inference costs (measured by token usage) alongside accuracy for various models and prompting strategies. Basic models like GPT-4o-mini and GPT-4o exhibit low accuracy and minimal cost, whereas advanced methods such as ToT Yao et al. (2023) and MACM Lei et al. (2024) improve accuracy at the expense of increased token consumption. Our method achieves a strong balance, achieving high accuracy with moderate token usage, outperforming MACM, which is more costly yet less accurate. Proprietary models like o1-mini and o1-preview attain slightly higher accuracy but incur the highest cost, underscoring their greater resource demands. Overall, Meta-Reasoner provides a cost-effective and scalable solution for reasoning-intensive tasks.
Effectiveness of Meta-reasoner.
Figure 3 demonstrates the cumulative rewards across iterations. We compare our MAB-based approach with a baseline that directly prompts an LLM to select an arm (or “strategy”), referred to as Baseline (Direct Arm Selection); the prompt details are in Figures 47. Results show that the MAB-based meta-reasoner (using LinUCB Li et al. (2012)) consistently outperforms both direct LLM decision-making and random search across two tasks (Game of 24 and TheoremQA) and two model scales (GPT-4o-mini and Gemini-Exp-1206). While direct LLM prompting yields reasonable initial performance and random search requires minimal setup, neither approach effectively balances exploration and exploitation. In contrast, the MAB updating strategy leverages feedback from prior iterations to adaptively refine action selection (e.g., choosing an appropriate strategy based on CoT reasoning), steadily increasing cumulative rewards.
6 Conclusion
In this work, we introduce Meta-Reasoner, a meta-reasoning framework designed to enhance the reasoning capabilities of LRMs and optimize the inference-time reasoning efficiency. By operating as an “advisor”, meta-reasoner dynamically evaluates the reasoning process and provides high-level strategic guidance, addressing key limitations of o1-like reasoning chains, such as compounding errors and inefficiency in inference computing. Unlike conventional reasoning approaches, Meta-Reasoner focuses on global oversight rather than granular step-by-step processes, enabling LRMs to avoid unproductive lines of thought and better allocate computational resources. The experiments highlight the potential of dynamic reasoning chains to overcome inherent challenges in the LLM reasoning process and also show promise in broader applications, offering a scalable and adaptable solution for reasoning-intensive tasks.
Limitations
Our proposed Meta-Reasoner framework, while effective at enhancing inference-time reasoning, remains limited to text-based problems and struggles to address tasks requiring other modalities, such as geometry. Overcoming these challenges calls for further advancements in the model’s cognitive capabilities.
Appendix A Baselines
We consider several established prompting methods as baselines as follows:
  • Chain-of-thought (CoT) Wei et al. (2022): A prompting technique that encourages models to generate intermediate reasoning steps to enhance problem-solving capabilities.
  • Self-Consistent Chain of Thought (SC-CoT) Wang et al. (2022): An extension of CoT that improves reasoning consistency by generating multiple reasoning chains and selecting the most consistent answer.
  • Multi-Chain Reasoning (MCR) Yoran et al. (2024): enhances SC-CoT by having another LLM to assess and integrate content among the sampled reasoning chains to generate the final consistent answer.
  • Tree of Thoughts (ToT) Yao et al. (2023): A method that explores multiple reasoning paths in a tree structure, allowing the model to consider various possibilities before arriving at a conclusion by tree search algorithms.
  • Reflexion Shinn et al. (2024): A framework that enables models to reflect on their reasoning process, iteratively refining their answers based on feedback.
  • MACM Lei et al. (2024): A multi-agent system to refine the reasoning based on iterative condition mining.
Appendix B Computational Efficiency
B.1 Why Measure Efficiency Through Token Usage?
Evaluating the computational efficiency of LLMs is a critical yet nuanced task. A natural metric is wall-clock time, which directly reflects the real-world latency experienced during LLM inference. However, when interacting with LLMs primarily through API calls, as what we do in this paper, wall-clock time measurements become unreliable and difficult to interpret. This unreliability arises from several uncontrollable factors, including network latency, server-side load balancing, and dynamic optimizations on the provider’s infrastructure. These external variables introduce noise and variability that obscure the intrinsic computational characteristics of the reasoning method itself.
To address these challenges, we adopt token usage as a more stable and conceptually meaningful proxy for computational cost. Token consumption directly correlates with the amount of computation performed by the LLM and is widely recognized in recent literature as a standard efficiency metric Liu et al. (2025); Li et al. (2025b). By focusing on token reduction, we provide a consistent basis for comparing reasoning strategies that is less susceptible to external system variability.
B.2 Iterative Reasoning and Memory Management
Unlike traditional CoT methods that generate reasoning process in a single pass, Meta-Reasoner pauses generation periodically to evaluate the reasoning progress and adaptively switch the reasoning strategies using a multi-armed bandit-based algorithm. This iterative design adds additional memory overhead due to maintaining context across multiple LLM calls. To mitigate this, we summarize the entire reasoning history into compact progress reports (§4.2), greatly reducing the token load per iteration. For example, a typical 1000-token CoT step may consume approximately 500MB of KV cache (e.g., using LLaMa-2-7B in float16 precision), while our method’s 5 iteration with 100-token steps peak at roughly 50MB each. Additionally, early pruning of unlikely reasoning paths further controls memory usage.
B.3 Managing Reward Model Overhead
To minimize the computational cost of the reward model, our design relies on summarized progress reports rather than the full reasoning history, thereby reducing token usage and associated processing overhead. The reward computation is further optimized by employing the LinUCB algorithm, a MAB method whose complexity scales linearly with the number of strategies. Since the number of strategies remains relatively small—typically between 3 and 5 in fixed settings and up to 10 to 15 in dynamic settings—the computational burden remains manageable. As demonstrated in Figure 2, the savings achieved by avoiding unproductive reasoning paths significantly outweigh the modest overhead introduced by the strategy selection process, resulting in an overall efficient reward computation framework.
Appendix C Stability of Dynamic Strategy Generation
Dynamically adding or refining strategies, as discussed in §4.3 (dynamic contextual bandit), enhances flexibility but risks instability. To address this trade-off, our approach incorporates several key mechanisms:
We begin with a small set of verified strategies (Table 1) to establish a stable foundation and prevent arbitrary generation. Building on this, the LLM generate contextually relevant and precise strategies, like “Pause to clarify and disambiguate the current reasoning” and “Simplify by decomposing the task into smaller, manageable sub-tasks”, demonstrate how the meta-reasoner effectively targets specific reasoning bottlenecks.
Second, we employ a contextual multi-armed bandit to balance exploration and exploitation, filtering out suboptimal strategies during LLM inference time. As shown in Table 6, the dynamic bandit variant achieves 89% accuracy on task Game-of-24, significantly surpassing fixed strategy sets (65%–72%). This highlights the dynamic bandit’s ability to prioritize effective strategies while discarding less useful ones.
Third, the reward function described in §4.3 provides immediate feedback on strategy performance, enabling rapid deprioritization of confusing or unproductive strategies. Figure 3 shows a consistent increase in cumulative rewards over iterations, indicating that the system maintains coherent reasoning as it incorporates new strategies.
Together, these mechanisms enable the dynamic contextual bandit to maintain stability without compromising adaptability. The observed performance gains on both the Game-of-24 and TheoremQA datasets (Table 6) show that the benefits of dynamic strategy generation substantially outweigh potential instability concerns in practice.
Algorithm 1 Meta-Reasoner: Meta-Reasoning with Contextual Multi-Armed Bandits
0:  LRM $M$, bandit $\mathcal{B}$, initial strategy set $\mathcal{A}_{1}$, maximum rounds $T$
0:  Final answer $A_{\mathrm{final}}$
1:  $C_{0}\leftarrow\emptyset$;  .$\mathcal{B}.\mathrm{Initialize}(\mathcal{A}_{1})$
2:  $G_{0}\leftarrow$ default strategy
3:  for $t=1$ to $T$ do
4:     if $t>1$ then
5:        $P_{t-1}\leftarrow f(C_{t-1})$ // Summarize the existing CoT
6:        $x_{t-1}\leftarrow\mathrm{FeatureExtract}(P_{t-1})$ // Extract features for context
7:        (Optional): $\mathcal{A}_{t}\leftarrow\mathcal{A}_{t-1}\cup\{\text{new strategies}\}$ // Update strategy set dynamically
8:        $a_{t-1}\leftarrow\arg\max_{a\in\mathcal{A}_{t}}\mathrm{Score}_{\mathcal{B}}(x_{t-1},a)$ // Select strategy using bandit
9:        $G_{t}\leftarrow a_{t-1}$ // Set current guidance
10:     else
11:        $G_{t}\leftarrow G_{0}$ // Use default guidance for the first iteration
12:     end if
13:     $s_{t}\leftarrow M(C_{t-1},G_{t})$ // Generate new CoT with integrated guidance
14:     $C_{t}\leftarrow C_{t-1}\cup\{s_{t}\}$ // Append new reasoning step to the CoT
15:     $r_{t}\leftarrow\mathrm{ComputeReward}(C_{t})$ // Compute reward based on the updated CoT
16:     if $t>1$ then
17:        .$\mathcal{B}.\mathrm{Update}(x_{t-1},a_{t-1},r_{t})$ // Update bandit with observed feedback
18:     end if
19:     if termination condition met then
20:        break
21:     end if
22:  end for
23:  $A_{\mathrm{final}}\leftarrow\mathrm{ExtractAnswer}(C_{t})$
24:  return  $A_{\mathrm{final}}$
Figure 4: Prompt Demonstration (Page-1)
Figure 5: Prompt Demonstration (Page-2)
Figure 6: Prompt Demonstration (Page-3)
Figure 7: Prompt Demonstration (Page-4)

### v4

@@
 Meta-Reasoner: Dynamic Guidance for Optimized Inference-time Reasoning in Large Language Models
-Abstract
-Large Language Models (LLMs) increasingly rely on prolonged reasoning chains to solve complex tasks.
-However, this trial-and-error approach often leads to high computational overhead and error propagation, where early mistakes can derail subsequent steps.
-To address these issues, we introduce Meta-Reasoner, a framework that dynamically optimizes inference-time reasoning by enabling LLMs to “think about how to think.” Drawing inspiration from human meta-cognition and dual-process theory, Meta-Reasoner operates as a strategic advisor, decoupling high-level guidance from step-by-step generation.
-It employs contextual multi-armed bandits to iteratively evaluate reasoning progress and select optimal strategies (e.g., backtrack, clarify ambiguity, restart from scratch, or propose alternative approaches), and reallocates computational resources toward the most promising paths.
-Our evaluations on mathematical reasoning and puzzles highlight the potential of dynamic reasoning chains to overcome inherent challenges in the LLM reasoning process and also show promise in broader applications, offering a scalable and adaptable solution for reasoning-intensive tasks.
-Meta-Reasoner: Dynamic Guidance for Optimized Inference-time
-Reasoning in Large Language Models
-Yuan Sui1, Yufei He1, Tri Cao1, Simeng Han2, Yulin Chen1, Bryan Hooi1 1 National University of Singapore 2 Yale University yuan.sui@u.nus.edu, bhooi@comp.nus.edu.sg
-1 Introduction
-Recent advances in o1-like and r1-like [^1] O1 (from OpenAI) and R1 (from Deepseek) are two advanced LLM models designed for reasoning.
-reasoning have enabled large language models (LLMs) to achieve remarkable performance on complex tasks such as mathematics Patel et al.
-(2024); Lightman et al.
-(2023), science Rein et al.
-(2023), and logical puzzles Lei et al.
-(2024); Yao et al.
-(2023).
-By simulating multi-step, human-like deliberation Yao et al.
-(2024), these methods allow LLMs to decompose problems into smaller subproblems, test hypotheses, reflect on intermediate results, and iteratively refine their solutions.
-This extended reasoning process enables systematic exploration of ideas, verification of partial conclusions, and progressive improvement before producing a final answer.
-Such capabilities are particularly valuable in domains demanding rigorous logical reasoning Chenghao Yang (2024).
-Despite these advances, o1/r1-like reasoning remains fundamentally challenged by its trial-and-error nature: models generate numerous candidate reasoning paths, discard flawed ones, and gradually converge on solutions.
-While this flexibility facilitates exploration of diverse strategies, it often incurs substantial computational overhead Snell et al.
-(2024); Manvi et al.
-(2024) and is vulnerable to error propagation, where early mistakes accumulate and compromise subsequent steps Lei et al.
-(2024); Yao et al.
-(2023); Gandhi et al.
-(2024).
-Some iterative methods incorporate partial revision or backtracking Gandhi et al.
-(2024); Li et al.
-(2025a), but these approaches tend to be ad-hoc and limited to correcting errors within a narrow reasoning window.
-Crucially, they lack a systematic mechanism to assess whether an entire reasoning trajectory remains promising or should be abandoned.
-As a result, LLMs risk becoming “stuck” on unproductive reasoning paths, wasting valuable computational resources without recognizing when a strategic pivot is necessary.
-A critical challenge, therefore, is to enable LLMs to manage their reasoning budget more effectively—prioritizing promising directions while adapting or discarding ineffective strategies during inference time.
-To address this challenge, we propose Meta-Reasoner, a specialized meta-reasoning module that operates alongside the LLM to enhance its reasoning capabilities.
-Acting as a high-level advisor, the meta-reasoner dynamically evaluates the reasoning process and provides strategic guidance or redirection when progress stalls.
-Unlike the LLM, which focuses on detailed stepwise generation, the meta-reasoner maintains a global perspective, assessing overall progress and strategy from a high level.
-Meta-Reasoner operates in iterative rounds: First, the LLM generates partial chain-of-thought reasoning chains and a concise “progress report” summarizing its current state.
-Meta-Reasoner then reviews this report and offers high-level feedback - such as restarting reasoning with a different approach, refining existing ideas, or focusing on specific subproblems.
+Abstract.
+Large Language Models (LLMs) struggle with high computational time and error propagation during inference time, especially for complex tasks like math, puzzles, or coding requiring multi-step thinking.
+While existing reasoning models with chain-of-thoughts (CoT) can enable LLMs to do step-wise analysis and reflection, they often face the issue of wasting computation on less productive solutions and fail to make progress during inference time.
+In this paper, we propose Meta-Reasoner, a new framework to enable LLMs “Think about how to think”, i.e., optimize the inference compute by adjusting strategies on how to reason during inference time.
+Inspired by dual-process theory, our method decouples the high-level strategy generation (e.g., backtracking, switching approaches, or restarting) from stepwise CoT generation via a lightweight progress report.
+The strategy module only consider the summarized version from the previous CoTs to propose new strategies accordingly.
+We employ the contextual multi-armed bandits (CMABs) for this module to iteratively evaluate the previous reasoning states and dynamically adjust the strategy to avoid reasoning get stuck in less productive paths during inference.
+Evaluations on math problems (e.g., Game-of-24, TheoremQA) and scientific problems (e.g., SciBench) demonstrate that our method improves performance by 9-12% over previous SOTA methods while reducing inference time by 28-35%.
+This approach also generalizes to other domains like creative writing, demonstrating its versatility for diverse reasoning-intensive problems using LLMs.
+Large Language Models, Meta-Learning, Dual-Process Theory, Inference Time Compute
+copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Natural language generation
+1. Introduction
+Recent models incentivize stepwise chain-of-thought (CoT) reasoning (e.g., o1 from OpenAI (OpenAI et al., 2024), and r1 from Deepseek (DeepSeek-AI et al., 2025)) have achieve remarkable performance on complex reasoning tasks, such as math (Patel et al., 2024; Lightman et al., 2023), science (Rein et al., 2024), and logical puzzles (Lei et al., 2024; Yao et al., 2023).
+By simulating human-like deliberation (Yao et al., 2024; Wei et al., 2022), these methods allow large language models (LLMs) to decompose problems into smaller subproblems, test hypotheses, reflect on intermediate results, or iteratively refine their solutions (Cao et al., 2024).
+This extended reasoning process enables LLMs to achieve progressive improvement during inference time before producing a final answer (Chenghao Yang, 2024).
+Despite these advances, these multi-step reasoning paradigm remains challenged by its trial-and-error nature: models generate numerous candidate reasoning paths, explore each of them and discard the error ones and gradually converge on final solution.
+While this facilitates exploration of diverse strategies, it suffers from two critical limitations: (1) computational inefficiency, where models waste resources exploring dead-end paths during inference (Snell et al., 2024; Manvi et al., 2025); and (2) error propagation, where early mistakes accumulate and compromise subsequent steps (Lei et al., 2024; Yao et al., 2023; Gandhi et al., 2024).
+While models like o1 can perform limited backtracking, such as reevaluating individual steps or self-verification to revise their outputs (Gandhi et al., 2024; Li et al., 2025).
+They still lack a systematic mechanism to holistically revise or redirect their reasoning.
+This inability to dynamically adapt their strategy can lead to the model becoming stuck on unproductive lines of thought or inefficiently exploring the problem space without recognizing a strategic pivot is necessary (Gao et al., 2024).
+Additionally, as LLMs are tackling more challenging problems which require longer reasoning paths, these limitations will be critical bottleneck for the LLM development.
+A critical challenge, therefore, is to enable LLMs to manage their reasoning budget more effectively, i.e., prioritizing promising directions while adapting or discarding ineffective strategies during inference time.
+Addressing this requires a novel approach to provide adaptive oversight from the reasoning process.
+In this paper, we propose Meta-Reasoner, a specialized module that operates alongside the LLM to dynamically optimize its reasoning strategies during inference time.
+Acting as a high-level advisor, Meta-Reasoner dynamically evaluates the reasoning process of the LLM and provides strategic guidance or redirection when progress stalls.
+Inspired by dual-process theory (Didolkar et al., 2024), Meta-Reasoner decouples high-level strategies from detailed stepwise CoT generation from LLM via a lightweight progress report module.
+The strategy generation module only consider the summarized version from the previous CoTs to propose new strategies accordingly.
+The system operates iteratively: (1) the LLM generates partial chain-of-thought reasoning steps (§4.1) and a summarizes the reasoning states into a lightweight progress report (§4.2).
+(2) Meta-Reasoner then reviews this report and using contextual multi-armed bandits (CMABs) (§4.3) to generate high-level strategies (e.g., backtrack to step 3, switch decomposition approach, or restart with new axioms).
+(3) The LLM then executes refined steps guided by this feedback.
 This setup allows Meta-Reasoner to concentrate on overall strategy rather than getting involved in the granular details of the LLM’s reasoning.
-Overall, Meta-Reasoner helps prevent the LLM from getting stuck or spending resources on unproductive lines of inquiry during the inference time.
-Overall, the main contributions are as follows:
-  • We propose a meta-reasoning module that provides high-level oversight and guidance to the LLM’s reasoning, helping to prevent it from getting stuck on unproductive solution paths during inference.
+Based on our experiments results (§5), we find that our proposed method demonstrate better performance compared to other baselines on complex tasks such as math, puzzles, or logic tasks, and the performance of our method can be generalized to some other domains like creative writing.
+Overall, the main contributions of this paper are as follows:
+  • We propose a meta-reasoning module that provides high-level strategy and guidance for LLM reasoning, to prevent it from getting stuck on unproductive solution paths during inference.
   • We design a lightweight progress reporting mechanism, enabling efficient communication of the LLM’s reasoning state to the meta-reasoner with minimal overhead.
-  • We empirically show that our Meta-Reasoner improves accuracy and efficiency on challenging math and scientific reasoning benchmarks (e.g., Game of 24, TheoremQA, and SciBench) compared to existing strong baselines.
+  • The framework requires no task-specific tuning and generalize to math, science, puzzles, and other creative problems require complex reasoning.
+  • We empirically show that our Meta-Reasoner improves accuracy and efficiency on challenging benchmarks (e.g., Game of 24, TheoremQA, and SciBench) compared to existing strong baselines.
 Our results highlight the framework’s potential as a new scalable solution to inference-time reasoning.
-2 Related Works
+2. Related Works
 In this section, we first review the challenges of complex reasoning in LLMs and the limitations of chain-of-thought (CoT) methods.
@@
 Finally, we draw inspiration from cognitive science, framing our approach within the meta-cognition and dual-process theory to motivate the design of our Meta-Reasoner framework.
-Complex Reasoning in LLMs
-The introduction of CoT reasoning has transformed how LLMs tackle complex problems by decomposing tasks into intermediate steps Lee et al.
-(2025).
-Recent models such as OpenAI’s o1 and o3, and Deepseek’s r1, have achieved state-of-the-art performance across diverse domains by leveraging CoT-like reasoning Manvi et al.
-(2024); Li et al.
-(2025a); Kudo et al.
-(2024); Sui et al.
-(2024b).
-However, the inherent sequential dependency in CoT limits robustness: errors in early steps can propagate and degrade overall performance Snell et al.
-(2024).
-Moreover, when confronted with complex reasoning tasks, CoT methods may become trapped in infinite reasoning loops Lee et al.
-(2025); Sui et al.
-(2024a, c).
+Complex Reasoning in LLMs.
+The introduction of CoT reasoning has transformed how LLMs tackle complex problems by decomposing tasks into intermediate steps (Lee et al., 2025).
+Recent models such as OpenAI’s o1 and o3, and Deepseek’s r1, have achieved state-of-the-art performance across diverse domains by leveraging CoT-like reasoning (Manvi et al., 2025; Li et al., 2025; Kudo et al., 2024; Sui et al., 2025a).
+However, the inherent sequential dependency in CoT limits robustness: errors in early steps can propagate and degrade overall performance (Snell et al., 2024).
+Moreover, when confronted with complex reasoning tasks, CoT methods may become trapped in infinite reasoning loops (Lee et al., 2025; Sui et al., 2025b).
 To address these challenges, we propose Meta-Reasoner, a meta-reasoner that monitors and adapts the reasoning strategy based on the progress of CoT.
 By integrating mechanisms such as backtracking and self-verification, Meta-Reasoner provides a holistic perspective that evaluates the overall reasoning trajectory to prevent stagnation and reduce wasted computation on unproductive lines of thought.
-Backtracking and Self-Verification
-To mitigate the limitations of CoT-like reasoning, recent methods have explored backtracking and self-verification techniques Yao et al.
-(2023); Besta et al.
-(2023); Gandhi et al.
-(2024).
+Backtracking and Self-Verification.
+To mitigate the limitations of CoT-like reasoning, recent methods have explored backtracking and self-verification techniques (Yao et al., 2023; Besta et al., 2023; Gandhi et al., 2024).
 For example, Weng et al.
@@
 (2023) propose generating multiple candidate reasoning chains alongside a verifier mechanism that identifies and backtracks on erroneous steps.
-These approaches extend beyond post-hoc validation by enabling dynamic strategy adjustments during inference Lightman et al.
-(2023), thereby limiting error propagation in lengthy reasoning chains and mitigating infinite reasoning loops.
-Building on these efforts, our Meta-Reasoner framework employs instructions to (1) restart from scratch with alternative strategies, (2) backtrack to the error point, and (3) continue with targeted suggestions.
-Further details on this strategy are provided in §4.3.
-Meta-Cognition & Dual-Process Systems
-From a cognitive science perspective, meta-cognition involves higher-order processes that allow individuals to monitor, evaluate, and adjust their cognitive strategies Gao et al.
-(2024); Yoran et al.
-(2024).
-This reflective thinking—often characterized as System 2 in dual-process theories Havrilla et al.
-(2024)—is vital for tasks requiring careful deliberation and error correction Didolkar et al.
-(2024).
+These approaches extend beyond post-hoc validation by enabling dynamic strategy adjustments during inference (Lightman et al., 2023), thereby limiting error propagation in lengthy reasoning chains and mitigating infinite reasoning loops.
+Building on these efforts, our Meta-Reasoner framework employs instructions such as (1) restart from scratch with alternative strategies, (2) backtrack to the error point, and (3) continue with targeted suggestions to optimize the reasoning process.
+Further details on this strategy design are provided in §4.3.
+Meta-Cognition & Dual-Process Systems.
+From a cognitive science perspective, meta-cognition involves higher-order processes that allow individuals to monitor, evaluate, and adjust their cognitive strategies (Gao et al., 2024; Yoran et al., 2023) during thinking process.
+This reflective thinking—often characterized as System 2 in dual-process theories (Havrilla et al., 2024)—is vital for tasks requiring careful deliberation and error correction (Didolkar et al., 2024).
 Drawing on these insights, our Meta-Reasoner framework can be viewed as analogous to dual-process systems: the LLM generates CoT steps akin to System 1, while the Meta-Reasoner provides high-level strategic oversight, analogous to System 2, guiding or redirecting reasoning as needed.
 This separation of responsibilities balances efficiency with robust problem-solving, allowing the LLM to handle routine inferences and the Meta-Reasoner to intervene for strategic adjustments.
-3 Preliminary
-In complex reasoning tasks, a key challenge is deciding which strategy to use from multiple valid options.
-This problem can be naturally framed as a contextual multi-armed bandit (MAB) problem, a well-studied framework for making decisions that balance trying new options (exploration) and using what is already known to work (exploitation).
+3. Preliminary
+In complex reasoning tasks, a key challenge is deciding the best strategy from several valid options.
+Considering solving a complex math problem: researchers may employ strategies like decomposition, abstraction, heuristic validation, boundary testing or any other methods at any given step.
+The critical question is which strategy to use and when.
+This decision making problem aligns well with contextual multi-armed bandits (CMABs) problem (Slivkins, 2019), a well-studied framework for making such choices.
+The framework is designed to balance exploration (trying new strategies to learn) and exploitation (using strategies known to be effective).
+Figure 1. Demonstration: Contextual Multi-armed Bandits.
+Intuitive Understanding of CMAB.
 Imagine an agent faced with several strategies (called arms), and at each step, it observes some information about the current situation (called the context).
 Based on this context, the agent picks one strategy to apply and then receives feedback (a reward) indicating how well that strategy performed.
-The goal is to choose strategies over time to maximize the total reward.
+The goal of the CMAB problem is to choose strategies over time to maximize the total reward.
+As shown in Figure 1, it illustrates a dynamic strategy selection process using the CMAB approach.
+The controller manages three arms which correspond to three strategies—Strategy 1, Strategy 2, and Strategy 3—each with an associated Strategy Value (Q-value) that reflects its effectiveness.
+Based on feedback ($r_{t_{b}}$) from problem-solving attempts, the controller updates the probability distribution $\pi_{t_{b}}$ (illustrated as the “strategy choice probability”), which determines the likelihood of selecting each strategy (e.g., Strategy 1 might have a 50% chance, Strategy 2 a 30% chance, and Strategy 3 a 20% chance).
+The sample process then picks a specific strategy $\alpha_{i}$ (illustrated as the “chosen strategy”) from this distribution to apply to the current problem.
+The resulting performance feedback is fed back into the controller, which adjusts the Strategy Values and refines $\pi_{t_{b}}$ for the next iteration, optimizing the strategy selection over time.
+The bar chart visualizes the current distribution of $\pi_{t_{b}}$ across the strategies.
+Formalization of CMABs.
 Formally, at each time step $t$, the agent observes a context vector $x_{t}$ describing the current state and selects an arm $s_{t}$ from a set of possible strategies $\mathcal{S}$.
@@
 The agent aims to maximize the cumulative reward over $T$ steps:
-$R(T)=\sum_{t=1}^{T}r(s_{t},x_{t}).$ (1)
-A popular algorithm for this problem is LinUCB Li et al.
-(2012), which assumes the expected reward is approximately a linear function of the context.
+(1) $R(T)=\sum_{t=1}^{T}r(s_{t},x_{t}).$
+A popular algorithm for this problem is LinUCB (Li et al., 2010), which assumes the expected reward is approximately a linear function of the context.
 Specifically, for each arm $s$, there is an unknown parameter vector $\theta_{s}$ such that:
-$\mathbb{E}[r(s,x_{t})]\approx x_{t}^{\top}\theta_{s}.$ (2)
+(2) $\mathbb{E}[r(s,x_{t})]\approx x_{t}^{\top}\theta_{s}.$
 LinUCB maintains an estimate $\hat{\theta}_{s}$ of this parameter and a measure of uncertainty about the estimate.
 At each step, it selects the arm that maximizes a combination of the estimated reward and an uncertainty bonus:
-$s_{t}=\arg\max_{s\in\mathcal{S}}\left[x_{t}^{\top}\hat{\theta}_{s}+c\sqrt{x_{t}^{\top}A_{s}^{-1}x_{t}}\right],$ (3)
+(3) $s_{t}=\arg\max_{s\in\mathcal{S}}\left[x_{t}^{\top}\hat{\theta}_{s}+c\sqrt{x_{t}^{\top}A_{s}^{-1}x_{t}}\right],$
 where $A_{s}$ is a matrix capturing past observations for arm $s$, and $c$ controls how much the agent favors exploring uncertain options.
@@
 By using context to guide its choices, LinUCB adapts its strategy selection to the current situation, which aligns with our goal in Meta-Reasoner: to dynamically choose the most effective reasoning strategy during inference time.
-4 Methods
-Motivated by the intuition that LLMs should concentrate their computational efforts on more promising reasoning paths during inference time, we explore two key research questions in this paper: (1) How can language models dynamically allocate resources during inference to optimize reasoning and planning?; (2) What architectural design enables an effective separation between the reasoning process within the LLM and the meta-level guidance that oversees it?
-To address these questions, we propose a novel framework, Meta-Reasoner, which endows LLMs with the ability to “think about how to think”.
-Our framework supervises the reasoning process and dynamically guides the model to focus on more promising reasoning trajectories during inference time.
-Furthermore, Meta-Reasoner mitigates the limitations of conventional sequential reasoning, which may get stuck in suboptimal paths.
-We propose a “high-order" reasoning mechanism to balance exploration and exploitation using a Multi-Armed Bandit (MAB) algorithm.
-Figure 1: An illustration of the Meta-Reasoner workflow.
-In each round, the LLM produces a new reasoning step to extend its CoT reasoning.
-The CoT is then summarized into a progress report, which provides context for the meta-reasoner.
-Then meta-reasoner uses a contextual multi-armed bandit (either using a fixed contextual bandit or dynamic contextual bandit) to choose a guidance strategy.
+We provide the details about how to leverage these CMABs approaches in our system in Section 4.3.
+Figure 2.
+Workflow of Meta-Reasoner.
+In each round, LLM produces a new reasoning step to extend its CoT reasoning (§4.1).
+The CoT is then summarized into a progress report, which provides context for the meta-reasoner (§4.2).
+Then meta-reasoner uses a contextual multi-armed bandit (either using a fixed contextual bandit or dynamic contextual bandit) to choose a guidance strategy (§4.3).
 The selected strategy then guides the next reasoning step generation, to enable strategic redirection, error correction, and resource optimization.
-A reward is then computed from the progress report and used to update the bandit algorithm.
+A reward is then computed from the progress report and used to update the bandit algorithm (§4.3.1).
 The process repeats until the task is complete or the maximum number of rounds is reached.
-The meta-reasoning framework operates iteratively as illustrated in Figure 1.
+Table 1.
+Demonstration of contextual bandit pairs (i.e., diagnosis of the progress report (context) and the corresponding strategy (bandit)) for guiding the LLM reasoning process.
+The bottom four rows are some of the unique strategies generated by Dynamic Contextual Bandits illustrated in §4.3.3.
+Diagnosis Strategy
+Progress is insufficient or the current strategy seems ineffective. Restart from scratch and propose alternative strategies.
+There are mistakes in intermediate steps. Backtrack to the point where the error occurred.
+The current approach is working well. Continue and provide specific suggestions for the next steps.
+Ambiguous or conflicting intermediate results are observed. Pause to clarify and disambiguate the current reasoning, then reconcile the discrepancies.
+The reasoning process appears overly complex or convoluted. Simplify by decomposing the task into smaller, manageable sub-tasks.
+Evidence of error propagation or low confidence in certain sub-components. Perform targeted verification on critical steps and focus on areas with low confidence.
+Repetitive or circular reasoning patterns are detected. Reset to a previously successful checkpoint and explore alternative solution paths.
+4. Methods
+Motivated by the intuition that LLMs should concentrate their computational efforts on more promising reasoning paths during inference time, we explore two key research questions in this paper:
+  • (1) How can language models dynamically allocate resources during inference to optimize reasoning and planning?
+  • (2) What architectural design enables an effective separation between the reasoning process within the LLM and the meta-level guidance that oversees it?
+To address these questions, we propose Meta-Reasoner, which endows LLMs with the ability to “think about how to think”.
+Our framework supervises the LLM’s reasoning process and dynamically guides the model to focus on more promising reasoning trajectories during inference time.
+The framework operates iteratively as illustrated in Figure 2.
 At each round $t$, the reasoning process comprises three steps: (1) CoT generation by the LLM, (2) Progress Reporting to summarize the reasoning progress so far (i.e., this is partly for efficiency, and partly to help the meta-reasoner focus on its main goal of “advising” rather than being distracted by the details in the CoT), and (3) Strategy Generation by the meta-reasoner to optimize subsequent steps.
@@
 The complete algorithm of Meta-Reasoner is appended in Algorithm 1.
-4.1 Chain-of-Thought (CoT) Generation
-In the first step, the LLM generates a reasoning step to extend its CoT reasoning based on the user query.
-Starting from its reasoning history $C_{t-1}$ and the guidance $G_{t-1}$ provided by the meta-reasoner in the previous round, the LLM $M$ produces a new reasoning step $s_{t}$.
-This step is appended to the current CoT, forming $C_{t}=C_{t-1}\cup\{s_{t}\}$.
-By maintaining the full reasoning trajectory at each round, the model establishes a coherent foundation for evaluation and further refinement.
-This approach resembles the long-term reasoning demonstrated in models like o1, which generate extended CoTs.
-However, this reasoning process often resembles “trial-and-error”, potentially incurring unnecessary inference costs on unproductive paths.
-Moreover, due to its sequential nature, the process is prone to becoming trapped in suboptimal solutions.
-4.2 Progress Reporting
-After updating the CoT, we summarize the reasoning history $C_{t}$ into a concise progress report $P_{t}$.
-This summary captures key aspects of the reasoning trajectory, including progress toward the task goal, reasoning consistency, and significant updates thus far.
-The summarization function $f$ abstracts the detailed CoT into a simpler, more focused representation.
-This step is designed to be both computationally efficient and informative, enabling the meta-reasoner to evaluate high-level progress without being overwhelmed by the granular details of every reasoning step.
-While this may be viewed as an engineering heuristic, we observe that including essential information in the prompt unlocks the LLM’s capacity for “higher-order” thinking.
-Consequently, the LLM tends to generate more insightful and critical strategies, which are particularly valuable for complex reasoning tasks.
-Diagnosis Strategy
-Progress is insufficient or the current strategy seems ineffective. Restart from scratch and propose alternative strategies.
-[…]
-There are mistakes in intermediate steps. Backtrack to the point where the error occurred.
-[…]
-The current approach is working well. Continue and provide specific suggestions for the next steps.
-Ambiguous or conflicting intermediate results are observed. Pause to clarify and disambiguate the current reasoning, then reconcile the discrepancies.
-[…]
-The reasoning process appears overly complex or convoluted. Simplify by decomposing the task into smaller, manageable sub-tasks.
-[…]
-Evidence of error propagation or low confidence in certain sub-components. Perform targeted verification on critical steps and focus on areas with low confidence.
-[…]
-Repetitive or circular reasoning patterns are detected. Reset to a previously successful checkpoint and explore alternative solution paths.
-Table 1: Demonstration: Contextual bandit pair (i.e., diagnosis of the progress report (context) and the corresponding strategy (bandit)) for guiding the LLM’s reasoning process.
-Marked rows are some of the unique strategies generated by Dynamic Contextual Bandits.
-4.3 Meta-reasoner Strategy Generation
+4.1. Chain-of-Thought (CoT) Generation
+In the first step, the LLM generates a reasoning step $s_{t}$ to extend its CoT based on the user query.
+This step builds on the previous reasoning history $C_{t-1}$ and incorporates strategic guidance $G_{t-1}$ from the meta-reasoner in the prior round.
+The new reasoning step is appended to the CoT, forming $C_{t}=C_{t-1}\cup\{s_{t}\}$.
+By maintaining the complete reasoning trajectory at each iteration, the model maintains a coherent foundation for evaluating the refining the progress of reasoning process.
+This step resembles the long-term reasoning seen in models like o1 and r1, which generate extended CoTs during inference time.
+The key difference is that our meta-reasoner model actively intervenes the reasoning process by providing strategical guidance $G_{t-1}$, enabling the model “think about how to think”.
+4.2. Progress Reporting
+After updating the CoT, we apply a summarization function $f:C_{t}\rightarrow P_{t}$ to distill the most recent $t$ steps of the CoT into a concise progress report $P_{t}$.
+This report encapsulates critical elements of the reasoning trajectory, including the degree of progress toward the task objective, the logical consistency of the reasoning steps, and significant milestones or updates achieved.
+The function $f(\cdot)$ is designed to be computationally efficient, producing a compact summary that enables the meta-reasoner to assess high-level progress without processing the granular details of each step in $C_{t}$.
+While this summarization is heuristic in nature, we observe that it may trigger LLMs’ capacity for higher-order reasoning by including only essential information in $P_{t}$, which informs the meta-reasoner’s strategic guidance $G_{t}$.
+Empirically, this approach fosters more insightful and adaptive strategies, as the meta-reasoner can focus on critical patterns rather than exhaustive step-by-step details.
+We provide the detailed experiments in §5.3.
+4.3. Meta-reasoner Strategy Generation
 In the next step, the meta-reasoner evaluates the progress report $P_{t}$ and selects an appropriate strategy $G_{t}$ for LLM reasoning (the complete procedure is detailed in Algorithm 1).
-We formulate the generation of strategy as a contextual MAB problem and consider two settings: (1) a fixed-strategy formulation, where the meta-reasoner selects from a predefined set of strategies using a contextual bandit algorithm; and (2) an advanced setting, in which the meta-reasoner itself is an LLM-based agent capable of introducing or refining strategies dynamically.
-In both cases, the meta-reasoner employs the partial-feedback principle of MABs to adaptively select strategies based on a reward function that evaluates the quality of reasoning progress after applying the chosen strategy.
-We demonstrate the contextual bandit pair (i.e., diagnosis of the progress report (i.e., context) and the corresponding strategy (i.e., bandit) in Table 1.
-Progress Evaluation.
-A central objective of our evaluation mechanism is to quantify how effectively the model’s current reasoning advances toward the task goal (e.g., solving a complex problem), while also monitoring computational cost to promote efficiency.
-Concretely, we implement a reward function that tracks both solution progress (e.g., partial correctness, adherence to constraints) and resource usage (e.g., number of reasoning steps).
-This evaluator can be any suitable mechanism, including LLM-based verification or external scoring scripts.
-In our implementation, we leverage an LLM as the evaluator (with prompts referenced in Figures 47) to assess reasoning progress.
-The evaluator outputs a cumulative score, which is then used to update the MAB algorithm.
-Fixed Contextual Bandit.
+We formulate the generation of strategy as a CMAB problem (as defined in §3) and explore two settings: (1) a fixed-strategy formulation, where the meta-reasoner selects from a predefined set of strategies using a contextual bandit algorithm; and (2) an advanced setting, where the meta-reasoner, implemented as an LLM-based agent, dynamically generates or refines strategies.
+In both settings, the meta-reasoner leverages the partial-feedback mechanism of MABs to adaptively select strategies based on a reward function that evaluates the quality of reasoning progress after applying the chosen strategy $G_{t}$.
+We demonstrate the contextual bandit pair (diagnosis of the progress report (i.e., context) and the corresponding strategy (i.e., bandit) in Table 1.
+Algorithm 1 Meta-Reasoning with Contextual Multi-Armed Bandits
+0:  LLM $M$, Bandits $\mathcal{B}$, Initial strategy set $\mathcal{A}_{1}$, Maximum rounds $T$
+0:  Final answer $A_{\mathrm{final}}$
+1:  $C_{0}\leftarrow\emptyset$;  $\mathcal{B}.\mathrm{Initialize}(\mathcal{A}_{1})$
+2:  $G_{0}\leftarrow$ default strategy
+3:  for $t=1$ to $T$ do
+4:   if $t>1$ then
+5:    $P_{t-1}\leftarrow f(C_{t-1})$
+6:    $x_{t-1}\leftarrow\mathrm{FeatureExtract}(P_{t-1})$
+7:    (Optional): $\mathcal{A}_{t}\leftarrow\mathcal{A}_{t-1}\cup\{\text{new strategies}\}$
+8:    $a_{t-1}\leftarrow\arg\max_{a\in\mathcal{A}_{t}}\mathrm{Score}_{\mathcal{B}}(x_{t-1},a)$
+9:    $G_{t}\leftarrow a_{t-1}$
+10:   else
+11:    $G_{t}\leftarrow G_{0}$
+12:   end if
+13:   $s_{t}\leftarrow M(C_{t-1},G_{t})$
+14:   $C_{t}\leftarrow C_{t-1}\cup\{s_{t}\}$
+15:   $r_{t}\leftarrow\mathrm{ComputeReward}(C_{t})$
+16:   if $t>1$ then
+17:    $\mathcal{B}.\mathrm{Update}(x_{t-1},a_{t-1},r_{t})$
+18:   end if
+19:   if termination condition met then
+20:    break
+21:   end if
+22:  end for
+23:  $A_{\mathrm{final}}\leftarrow\mathrm{ExtractAnswer}(C_{t})$
+24:  return  $A_{\mathrm{final}}$
+4.3.1. Reward Modeling for Progress Report.
+The primary goal of our evaluation mechanism is to quantify how effectively the model’s current reasoning advances toward the task objective (e.g., solving a complex problem), while also monitoring computational cost to promote efficiency.
+We deine a reward function $R:R:P_{t}\times G_{t}\to\mathbb{R}$ that integrates two key components: (1) solution progress $S_{p}$, measuring correctness and adherence to query’s constraints (e.g., $S_{p}=w_{1}\cdot C_{c}+w_{2}\cdot C_{a}$, where $C_{c}$ is correctness, $C_{a}$ is adherence, and $w_{1},w_{2}$ are weights); and (2) resource usage $R_{u}$, reflecting computational cost (e.g., $R_{u}=-\alpha\cdot N_{s}$, where $N_{s}$ is the number of reasoning steps and $\alpha$ is a cost coefficient).
+The total reward is computed as $R(G_{t},P_{t})=\beta\cdot S_{p}+(1-\beta)\cdot R_{u}$, with $\beta\in[0,1]$ balancing the trade-off.
+This evaluation can be implemented via an LLM-based assessor (prompts detailed in Figures 69) or external scoring scripts, producing a cumulative score to update the CMAB algorithm.
+For our experiments, we use the same model as both CoT generation and reward modeling.
+4.3.2. Fixed Contextual Bandit.
 In the basic version of our framework, the meta-reasoner is modeled as a single contextual bandit that selects from a fixed, finite set of $K$ strategies.
-These strategies include instructions such as “continue and provide specific suggestions”, “restart from scratch”, “backtrack to the point where the error occurred”, or “propose alternative methods or perspectives to consider”, as detailed in Table 1 (rows without marked).
-At each round, the LLM produces a progress report summarizing its partial reasoning, the meta-reasoner transforms this progress report into a feature vector $x_{t}$ using a language model and applies a contextual bandit algorithm (e.g., LinUCB Li et al.
-(2012)) to select the best next strategy $a_{t}$.
+These strategies include instructions such as “continue and provide specific suggestions”, “restart from scratch”, “backtrack to the point where the error occurred”, or “propose alternative methods or perspectives to consider”, as detailed in Table 1.
+At each round, the LLM produces a progress report $P_{t}$ summarizing its recent several reasoning trajectory, the meta-reasoner then transforms this progress report into a feature vector $x_{t}$ using a language model and applies a contextual bandit algorithm (e.g., LinUCB (Li et al., 2010)) to select the best next strategy $a_{t}$.
 The LLM then executes that strategy and we collect the reward $r_{t}$ for $a_{t}$ based on the reward function.
 Through iterative MAB algorithm updating, the MAB algorithm learns to select appropriate strategies conditioned on the recent progress report.
-Dynamic Contextual Bandit.
-The fixed-arm formulation assumes a static set of strategies.
+4.3.3. Dynamic Contextual Bandit.
+The fixed-arm formulation assumes a static set of strategies $G_{0}$.
 In practice, the meta-reasoner may itself be an LLM capable of inventing new strategies over time.
-To accommodate dynamic strategies, we allow the meta-reasoner to propose or refine new strategies at round $t$, which generates an expanding collection of strategies, $G_{1}\subseteq\cdots\subseteq G_{t}$.
-Each newly proposed strategy becomes an arm in the contextual bandit.
-To encourage at least some exploration on this new arm, we initialize each arm with a blank or weak prior in the bandit’s parameters.
-We further analyze the stability of these new generated dynamic contextual bandits in Appendix §C.
-By explicitly separating low-level content generation (handled by the LLM) from high-level strategy decisions (governed by the meta-reasoner’s bandit), the system can effectively avoid getting stuck or wasting excessive resources on poor solution paths.
+To accommodate dynamic strategies, we allow the meta-reasoner to propose or refine new strategies at round $t$, which generates an expanding collection of strategies, $G_{1}\subseteq G_{2}\subseteq\cdots\subseteq G_{t}$.
+Each newly introduced strategy becomes an additional arm in the contextual multi-armed bandit framework.
+To encourage at least some exploration on this new arm, we initialize their bandit parameters with a neutral prior, such as $Q(a)=0$ for arm $a$ to avoid strong biases.
+We further analyze the stability of these new generated dynamic contextual bandits in §5.4.5, ensuring that the system avoids incorporating random or ineffective strategies by grounding new arms in contextual relevance and iterative feedback.
+By explicitly separating low-level content generation (handled by the LLM) from high-level strategy decisions (governed by the meta-reasoner’s bandit), the system can effectively avoid getting stuck or wasting excessive resources on poor reasoning paths.
 In domains where a predefined set of strategies is sufficient, the fixed-arm formulation can simplify the method deployment.
 While in more open-ended domains where novel tactics may emerge, dynamic-arm extensions give meta-reasoner more flexibility to evolve.
-5 Experiments
-In this section, we first introduce the experiment settings including datasets, backbone models and training details of MAB.
+5. Experiments
+In this section, we first introduce the experiment settings and training details of our method.
 We then present the main results of Meta-Reasoner with analysis regarding efficiency, rewards accumulation, and qualitative assessment of meta-reasoner output.
-We present the detailed baselines in Appendix A.
-5.1 Experiments Setup
+5.1. Experiments Setup
 Datasets.
-We test our method on several challenging datasets that demand complex reasoning and often involve lengthy thinking processes for the correct solutions.
-These includes (1) 24-point game Yao et al.
-(2023); (2) college-level scientific problem from SciBench Wang et al.
-(2024) and (3) math questions based on theorems from TheoremQA Chen et al.
-(2023).
+We mainly focus on tasks that demand complex reasoning and often involve lengthy thinking processes for the correct solutions.
+These includes (1) 24-point game (Yao et al., 2023); (2) college-level scientific problem from SciBench (Wang et al., 2023) and (3) math questions based on theorems from TheoremQA (Chen et al., 2023).
 For SciBench, we focus only on the math-related subsets (i.e., diff, stat, and calc).
@@
 We collect the training data for each task: (1) for the 24-point game, we random sample 50 queries from 4nums.com specifically excluding problems in ranks 901-1000 which were reserved for testing; (2) for TheoremQA, we randomly sample 30 mathematical reasoning queries from the dataset; (3) for SciBench, we randomly sample 30 queries from differential subsets including diff, stat, and calc from the entire dataset.
-These samples were used to iteratively update the LinUCB parameters for both fixed ($K$=3 or $K$=5 strategies) and dynamic strategy settings.
-We configure the training process using deterministic generation ($n$=1, Top_k=1, temperature=0) with specific max_token limits for CoT generation (512), meta-reasoner feedback (256), progress reports (512), and reward model outputs (4).
-The LinUCB exploration parameter was set to $c$=0.2.
-Experiments were ran for a maximum of $T$=30 iterations for the 24-point game, scibench, and $T$=100 for TheoremQA, with MAB parameters updated after each iteration based on a reward function that weighted objective completion (40%), progress quality (30%), efficiency (15%), and strategy alignment (15%).
-The reward function prompt is detailed in Figure 4-7.
+These samples were used to iteratively update the LinUCB parameters for both fixed (KK=3 or KK=5 strategies) and dynamic strategy settings.
+We configure the training process using deterministic generation (nn=1, Top_k=1, temperature=0) with specific max_token limits for CoT generation (512), meta-reasoner feedback (256), progress reports (512), and reward model outputs (4).
+The LinUCB exploration parameter was set to cc=0.2.
+Experiments were ran for a maximum of TT=30 iterations for the 24-point game, scibench, and TT=100 for TheoremQA, with MAB parameters updated after each iteration based on a reward function that weighted objective completion (40%), progress quality (30%), efficiency (15%), and strategy alignment (15%).
+The reward function prompt is detailed in Figure 6-9.
+Baselines.
+We consider several established prompting methods as baselines as follows: (1) Chain-of-thought (CoT) (Wei et al., 2022): A prompting technique that encourages models to generate intermediate reasoning steps to enhance problem-solving capabilities.
+(2) Self-Consistent Chain of Thought (SC-CoT) (Wang et al., 2022): An extension of CoT that improves reasoning consistency by generating multiple reasoning chains and selecting the most consistent answer.
+(3) Multi-Chain Reasoning (MCR) (Yoran et al., 2023): enhances SC-CoT by having another LLM to assess and integrate content among the sampled reasoning chains to generate the final consistent answer.
+(4) Tree of Thoughts (ToT) (Yao et al., 2023): A method that explores multiple reasoning paths in a tree structure, allowing the model to consider various possibilities before arriving at a conclusion by tree search algorithms.
+(5) Reflexion (Shinn et al., 2023): A framework that enables models to reflect on their reasoning process, iteratively refining their answers based on feedback.
+(6) MACM (Lei et al., 2024): A multi-agent system to refine the reasoning based on iterative condition mining.
+Backbone Models.
+We consider both LLMs and the recent Large Reasoning Models (LRMs) for our experiments.
+For the LLMs, we consider the closed-source models like gpt-4o, gpt-4o-mini (between Nov 2025 to Jan 2025) from OpenAI, and open-sourced models like meta-llama-3.1-8B-instruct from Meta, qwen-3-8B from Alibaba, phi-4 from Microsoft, gemini-experimental-1206 from Google, ds-r1-distill-llama-8B, ds-r1-distill-qwen-7B from Deepseek.
+For the LRMs, we consider the closed-source models like o1, o1-mini (In case we cannot break down the generation of o1 models through APIs, we cannot properly inject our meta-reasoner with o1-series models; we only provide the IO results for references).
+For the feature extraction mentioned in §4.3, we use text-embedding-3-small from OpenAI as the embedding model.
+To ensure the reproducibility of the experiments, we set $\mathrm{temperature}=0.7$ and $\mathrm{top\_p}=1.0$ for all models.
+We use the API service from OpenAI [^1] https://openai.com/ and OpenRouter [^2] https://openrouter.ai/ for our experiments which host detailed snapshots of the utilized model versions.
+5.2. Main Results
+We compare the accuracy of different prompting methods across different backbone models on SciBench (as shown in Table 4), 24-points game (as shown in Table 2) and TheoremQA (as shown in Table 3).
+We found that basic prompting strategies, such as CoT and SC-CoT, show limited effectiveness, achieving only 4% and 9% accuracy on 24-point games, respectively.
+Incorporating IO strategy with “Best of 100” samples improves accuracy to 33%, but it remains far behind advanced methods.
+Strategies like ToT illustrate the importance of exploring broader reasoning paths, with accuracy increasing from 45% to 74% as the search breadth expands from 1 to 5.
+Advanced iterative methods, such as Reflexion (53%) and MACM (80%), further demonstrate the value of refined reasoning frameworks.
+Our proposed Meta-Reasoner outperforms these approaches, achieving 89% accuracy with GPT-4o-mini and 92% with GPT-4o, showcasing its ability to dynamically guide reasoning, correct errors, and focus resources effectively.
+Compared to specialized models like o1-mini, our method equipped with much cheaper and generalized models like GPT-4o-mini delivers comparable performance, demonstrating its adaptability and scalability.
+Table 2. Performance of different models on 24-points Game (Yao et al., 2023) ($b$: search breadth)
+Method Accuracy (%)
+GPT-4o-mini + CoT (Yao et al., 2023) 4
+GPT-4o-mini + SC-CoT (Yao et al., 2023) 9
+GPT-4o-mini + IO (best of 100) (Yao et al., 2023) 33
+GPT-4o-mini + CoT (best of 100) (Yao et al., 2023) 49
+Gemini-Exp-1206 + IO (best of 100) (Yao et al., 2023) 38
+Gemini-Exp-1206 + CoT (best of 100) (Yao et al., 2023) 60
+GPT-4o-mini + ToT ($b=1$) (Yao et al., 2023) 32
+GPT-4o-mini+ ToT ($b=5$) (Yao et al., 2023) 65
+GPT-4o-mini + Reflexion (Shinn et al., 2023) 53
+GPT-4o-mini + MACM (Lei et al., 2024) 80
+GPT-4o-mini + Meta-Reasoner (our work) 89
+GPT-4o + Meta-Reasoner (our work) 92
+Gemini-Exp-1206 + Meta-Reasoner (our work) 94
+o1-mini + IO 89
+o1-preview + IO 93
+Table 3. Performance of different models on TheoremQA (Chen et al., 2023)
+Method Accuracy (%)
+GPT-4o-mini + CoT 39.46
+Gemini-Exp-1206 + CoT 43.12
+GPT-4o-mini + Reflexion (Shinn et al., 2023) 74.32
+GPT-4 Turbo + MACM (Lei et al., 2024) 79.41
+GPT-4o-mini + Meta-Reasoner (our work) 84.13
+Gemini-Exp-1206 + Meta-Reasoner (our work) 86.32
+Overall, the Meta-Reasoner framework provides a compatible approach to improving reasoning-intensive tasks, combining high accuracy with dynamic and efficient problem-solving strategies.
+The results on SciBench and TheoremQA also demonstrate similar findings and show that Meta-Reasoner generally achieves better performance compared to the baselines and the results are consistent across different models.
+Table 4. Performance of different models on SciBench Math Subsets (Wang et al., 2023)
 Method Diff(%) Stat(%) Calc(%)
@@
 GPT-4o-mini + MCR 40.12 58.21 67.42
-[…]
-GPT-4o-mini + MACM Lei et al. (2024) 54.78 67.13 65.77
-GPT-4o + MACM Lei et al. (2024) 61.42 78.32 76.72
+GPT-4o-mini + MACM (Lei et al., 2024) 54.78 67.13 65.77
+GPT-4o + MACM (Lei et al., 2024) 61.42 78.32 76.72
 GPT-4o-mini + Meta-Reasoner (our work) 60.32 73.64 80.23
 GPT-4o + Meta-Reasoner (our work) 67.14 83.29 84.17
-Table 2: Accuracy (%) comparison of different methods on the math-related subset of the SciBench dataset. Each column refers to the problem subset defined in Wang et al. (2024).
-Method Accuracy (%)
-GPT-4o-mini + CoT Yao et al. (2023) 4
-GPT-4o-mini + SC-CoT Yao et al. (2023) 9
-GPT-4o-mini + IO (best of 100) Yao et al. (2023) 33
-GPT-4o-mini + CoT (best of 100) Yao et al. (2023) 49
-Gemini-Exp-1206 + IO (best of 100) Yao et al. (2023) 38
-Gemini-Exp-1206 + CoT (best of 100) Yao et al. (2023) 60
-[…]
-GPT-4o-mini + ToT ($b=1$Yao et al. (2023) 32
-GPT-4o-mini+ ToT ($b=5$Yao et al. (2023) 65
-GPT-4o-mini + Reflexion Shinn et al. (2024) 53
-GPT-4o-mini + MACM Lei et al. (2024) 80
-GPT-4o-mini + Meta-Reasoner (our work) 89
-GPT-4o + Meta-Reasoner (our work) 92
-Gemini-Exp-1206 + Meta-Reasoner (our work) 94
-[…]
-o1-mini + IO 89
-o1-preview + IO 93
-Table 3: Accuracy(%) comparison of different prompting methods on 24-points game. $b$: Search breadth.
-Backbone Models.
-We consider both LLMs and the recent Large Reasoning Models (LRMs) for our experiments.
-For the LLMs, we consider the closed-source models like gpt-4o, gpt-4o-mini (between Nov 2025 to Jan 2025) from OpenAI, and open-sourced models like meta-llama-3.1-8B-instruct from Meta, phi-4 from Microsoft and gemini-experimental-1206 from Google.
-For the LRMs, we consider the closed-source models like o1, o1-mini (In case we cannot break down the generation of o1 models through APIs, we cannot properly inject our meta-reasoner with o1-series models; we only provide the IO results for references).
-For the feature extraction mentioned in §4.3, we use text-embedding-3-small from OpenAI as the embedding model.
-To ensure the reproducibility of the experiments, we set $\mathrm{temperature}=0.7$ and $\mathrm{top\_p}=1.0$ for all models.
-We use the API service from OpenAI [^2] https://openai.com/ and OpenRouter [^3] https://openrouter.ai/ for our experiments which host detailed snapshots of the utilized model versions.
-5.2 Main Results
-We compare the accuracy of different prompting methods across different backbone models on SciBench (as shown in Table 2), 24-points game (as shown in Table 3) and TheoremQA (as shown in Table 4).
-We found that basic prompting strategies, such as CoT and SC-CoT, show limited effectiveness, achieving only 4% and 9% accuracy on 24-point games, respectively.
-Incorporating IO strategy with “Best of 100” samples improves accuracy to 33%, but it remains far behind advanced methods.
-Strategies like ToT illustrate the importance of exploring broader reasoning paths, with accuracy increasing from 45% to 74% as the search breadth expands from 1 to 5.
-Advanced iterative methods, such as Reflexion (53%) and MACM (80%), further demonstrate the value of refined reasoning frameworks.
-Our proposed Meta-Reasoner outperforms these approaches, achieving 89% accuracy with GPT-4o-mini and 92% with GPT-4o, showcasing its ability to dynamically guide reasoning, correct errors, and focus resources effectively.
-Compared to specialized models like o1-mini, our method equipped with much cheaper and generalized models like GPT-4o-mini delivers comparable performance, demonstrating its adaptability and scalability.
-Overall, the Meta-Reasoner framework provides a compatible approach to improving reasoning-intensive tasks, combining high accuracy with dynamic and efficient problem-solving strategies.
-The results on SciBench and TheoremQA also demonstrate similar findings and show that Meta-Reasoner generally achieves better performance compared to the baselines and the results are consistent across different models.
-5.3 Ablation Study
-Method Accuracy (%)
-GPT-4o-mini + CoT 39.46
-Gemini-Exp-1206 + CoT 43.12
-[…]
-GPT-4o-mini + Reflexion Shinn et al. (2024) 74.32
-GPT-4 Turbo + MACM Lei et al. (2024) 79.41
-GPT-4o-mini + Meta-Reasoner (our work) 84.13
-Gemini-Exp-1206 + Meta-Reasoner (our work) 86.32
-Table 4: Accuracy(%) comparison of different prompting methods on TheoremQA Chen et al. (2023).
-Model Variant Game-of-24(%) TheoremQA(%)
+5.3. Ablation Study
+We further conduct an ablation study to analyze each component contribution of Meta-Reasoner.
+In specific, we consider the following setup: (1) w/o progress report: we replace the progress reporting process with directly considering the entire CoT history without summarization; (2) w/o MAB: instead of using MAB to select the proper strategy, we directly leverage an LLM to the decision making to provide the proper strategy for LRM reasoning.
+Table 5. Ablation study of Meta-Reasoner. Direct arm selection refers to prompting LLM to directly select a strategy based on recent progress report.
+Model Variant Game-of-24 TheoremQA
 GPT-4o-mini Full Method 89 84.13
@@
 w/o MAB (CoT) 11 43.12
-Table 5: Ablation study of Meta-Reasoner. Direct arm selection refers to prompting LLM to directly select a strategy based on recent progress report.
-Bandit Type Game-of-24(%) #US TheoremQA(%) #US
-Fixed (K=3) 65 3 72.34 3
-Fixed (K=5) 72 5 79.17 5
-Dynamic 89 14 84.13 21
-Table 6: Fixed vs. Dynamic Bandit Variants over GPT-4o-mini. #US: Number of Unique Strategies.
-Figure 2: We normalize the inference costs (based on token usage) in a range from 0 to 1 for easier comparison.
-We use gpt-4o-mini as the backend model for all the methods in this figure.
-For each method, key hyper-parameters (e.g., ${N}$ in Best of N, or tree size in ToT) are tuned to yield a baseline (lower point) and an extended (upper point) configuration, with dashed lines connecting these bounds.
-In this section, we conduct an ablation study to analyze each component contribution of Meta-Reasoner.
-In specific, we consider the following setup: (1) w/o progress report: we replace the progress reporting process with directly considering the entire CoT history without summarization; (2) w/o MAB: instead of using MAB to select the proper strategy, we directly leverage an LLM to the decision making to provide the proper strategy for LRM reasoning.
 In Table 5, we show that when removing progress reporting (“w/o Progress Report”), the overall performance moderately degrades and we hypothesize it is due to the concise intermediate summarizations can help the Meta-reasoner only consider the high-level strategy instead of being confused with too much details of the reasoning process.
@@
 The results highlight the benefit of flexibly allocating diverse reasoning strategies using LLM in-context learning capabilities.
-5.4 Analysis
-Figure 3: Cumulative reward of different settings across iteration.
-We compare our method using LinUCB with baseline (direct arm selection), and random search methods across two tasks—Game of 24 (top row) and TheoremQA (bottom row) using GPT-40-mini (left) and Gemini-Exp-1206 (right).
-Inference Efficiency.
-A common way to measure computational efficiency is wall-clock time, as it directly reflects real-world latency during inference.
-However, in our API-based setup, wall-clock time is unreliable due to external factors such as network latency and server load, which introduce noise unrelated to the model’s intrinsic computational cost.
-Therefore, we use token usage as a more stable and meaningful proxy, since it directly correlates with the amount of computation performed and is widely adopted in recent works Liu et al.
-(2025); Li et al.
-(2025b).
-The detailed instruction can be found in Appendix B.1.
-Figure 2 presents normalized inference costs (measured by token usage) alongside accuracy for various models and prompting strategies.
-Basic models like GPT-4o-mini and GPT-4o exhibit low accuracy and minimal cost, whereas advanced methods such as ToT Yao et al.
-(2023) and MACM Lei et al.
-(2024) improve accuracy at the expense of increased token consumption.
-Our method achieves a strong balance, achieving high accuracy with moderate token usage, outperforming MACM, which is more costly yet less accurate.
-Proprietary models like o1-mini and o1-preview attain slightly higher accuracy but incur the highest cost, underscoring their greater resource demands.
-Overall, Meta-Reasoner provides a cost-effective and scalable solution for reasoning-intensive tasks.
-Effectiveness of Meta-reasoner.
+Figure 3.
+Cumulative reward of different settings across iteration.
+We compare our method using LinUCB with baseline (direct arm selection), and random search methods across two tasks—Game of 24 (top row) and TheoremQA (bottom row) using GPT-4o-mini (left) and Gemini-Exp-1206 (right).
+(a) Game of 24 Task.
+(b) TheoremQA Task
+Figure 4. Inference time heatmap across different methods.
+5.4. Analysis
+5.4.1. Effectiveness of Meta-reasoner.
 Figure 3 demonstrates the cumulative rewards across iterations.
 We compare our MAB-based approach with a baseline that directly prompts an LLM to select an arm (or “strategy”), referred to as Baseline (Direct Arm Selection); the prompt details are in Figures 69.
-Results show that the MAB-based meta-reasoner (using LinUCB Li et al.
-(2012)) consistently outperforms both direct LLM decision-making and random search across two tasks (Game of 24 and TheoremQA) and two model scales (GPT-4o-mini and Gemini-Exp-1206).
+Results show that the MAB-based meta-reasoner (using LinUCB (Li et al., 2010)) consistently outperforms both direct LLM decision-making and random search across two tasks (Game of 24 and TheoremQA) and two model scales (GPT-4o-mini and Gemini-Exp-1206).
 While direct LLM prompting yields reasonable initial performance and random search requires minimal setup, neither approach effectively balances exploration and exploitation.
 In contrast, the MAB updating strategy leverages feedback from prior iterations to adaptively refine action selection (e.g., choosing an appropriate strategy based on CoT reasoning), steadily increasing cumulative rewards.
-6 Conclusion
+Table 6. Fixed vs. Dynamic Bandit Variants over GPT-4o-mini. #US: Number of Unique Strategies.
+Bandit Type Game-of-24(%) #US TheoremQA(%) #US
+Fixed (K=3) 65 3 72.34 3
+Fixed (K=5) 72 5 79.17 5
+Dynamic 89 14 84.13 21
+5.4.2. Inference Efficiency.
+We measure the inference efficieny of our proposed method.
+In Figure 4, we calculate the average inference time on different tasks across different models and methods.
+It demonstrates Meta-Reasoner’s significant advantages over existing reasoning frameworks across multiple benchmarks.
+Compared to o1-preview in the zero-shot setting, our method achieved roughly 51-55% inference time reduction while maintaining comparable performance.
+Among reasoning methods using identical base models, Meta-Reasoner consistently outperformed alternatives: requiring 29-33% fewer inference time than ToT (Yao et al., 2023), Reflexion (Shinn et al., 2023), MACM (Lei et al., 2024) and Best-of-N techniques.We further provide the demonstration of the token usage comparison in Figure 5.
+The efficiency results positions Meta-Reasoner as a scalable solution for complex reasoning tasks where computational resources are constrained.
+(a) Token Usage on Game of 24 Task.
+(b) Token Usage on TheoremQA Task
+Figure 5. Token Usage Across Different Methods on Game-of-24 (Yao et al., 2023) and TheoremQA (Chen et al., 2023) Tasks.
+5.4.3. Iterative Reasoning and Memory Management.
+Unlike traditional CoT methods that generate reasoning process in a single pass, Meta-Reasoner pauses generation periodically to evaluate the reasoning progress and adaptively switch the reasoning strategies using a multi-armed bandit-based algorithm.
+This iterative design adds additional memory overhead due to maintaining context across multiple LLM calls.
+To mitigate this, we summarize the entire reasoning history into compact progress reports (§4.2), greatly reducing the token load per iteration.
+As demonstrated in Figure 4, the savings achieved by avoiding unproductive reasoning paths significantly outweigh the modest overhead introduced by the strategy selection process, resulting in an overall efficient reward computation framework.
+5.4.4. Generalizability of Meta-Reasoner.
+To analyze the generalizability of our proposed method, we conduct additional experiment on creative writing, a domain that is quite different from math or logical reasoning.
+Following the instructions from ToT (Yao et al., 2023), we test our method on creative writing tasks where the input consists of four random sentences, and the output is a passage of four paragraphs ending with those sentences.
+We used the same evaluation metrics as ToT, which involves a GPT-4 zero-shot prompt providing a 1-10 scalar coherency score.
+Table 7. Performance on Creative Writing Task (Yao et al., 2023).
+Method GPT-4 Coherency Score
+IO (Input-Output) 6.19
+IO + Prompt Refined (k=5) 7.67
+CoT (Chain of Thought) (Wei et al., 2022) 6.93
+ToT (Tree of Thoughts) (Yao et al., 2023) 7.56
+ToT + Prompt Refined (Yao et al., 2023) 7.91
+Meta-Reasoner 7.68
+The results are shown in Table 7.
+Meta-Reasoner achieves a coherency score of 7.68, which is competitive with specialized methods like ToT + Refine (7.91) and significantly outperforms basic approaches such as IO (6.19) and CoT (6.93).
+These results indicate that our framework effectively transfers reasoning capabilities to creative tasks where objectives are less precisely defined, thereby supporting the claim that our method is indeed domain-agnostic.
+5.4.5. Stability of Dynamic Strategy Generation.
+Dynamically expanding or refining the strategy set, as outlined in Section 4.3, enhances adaptability but poses risks of instability due to potential random strategy generation.
+To mitigate this, our method incorporates three stabilizing mechanisms:
+(1) Initial Stable Foundation: We start with a curated set of verified strategies (e.g., Table 1), such as $g_{1}=$ ”Pause to clarify and disambiguate reasoning” and $g_{2}=$ ”Decompose task into sub-tasks,” ensuring a robust baseline.
+The LLM generates new strategies $g_{t}$ based on the current progress report $P_{t}$, constrained by contextual relevance, preventing arbitrary additions.
+(2) Exploration-Exploitation Balance: The CMAB employs an $\epsilon$-greedy policy to balance exploration and exploitation, with the probability of selecting a new arm ($a_{t}\in G_{t}\setminus G_{t-1}$) governed by $\epsilon_{t}=\frac{1}{t}$.
+Suboptimal strategies are filtered out during inference based on their empirical reward $R_{t}(a_{t},P_{t})$.
+For instance, on the Game-of-24 task, the dynamic bandit achieves 89% accuracy, outperforming fixed-strategy sets (65%–72%), as shown in Table 6, demonstrating effective prioritization of viable strategies.
+(3) Reward-Driven Feedback: The reward function $R:G_{t}\times P_{t}\to[0,1]$ (detailed in Section 4.3) provides real-time evaluation of strategy efficacy.
+Strategies yielding low rewards (e.g., $R_{t}<0.3$)) are deprioritized, ensuring rapid elimination of ineffective options.
+Figure 3 illustrates a monotonic increase in cumulative reward $\sum_{i=1}^{t}R_{i}$, reflecting stable reasoning as new strategies are integrated.
+Collectively, these mechanisms ensure the dynamic CMAB remains stable by anchoring strategy generation in context, filtering via bandit optimization, and refining through reward feedback.
+Empirical results on the Game-of-24 and TheoremQA datasets (Table 6) confirm that the adaptive benefits of dynamic strategy generation outweigh instability risks, with performance gains of 17%–24% over static approaches.
+6. Conclusion
 In this work, we introduce Meta-Reasoner, a meta-reasoning framework designed to enhance the reasoning capabilities of LRMs and optimize the inference-time reasoning efficiency.
@@
 The experiments highlight the potential of dynamic reasoning chains to overcome inherent challenges in the LLM reasoning process and also show promise in broader applications, offering a scalable and adaptable solution for reasoning-intensive tasks.
-Limitations
-Our proposed Meta-Reasoner framework, while effective at enhancing inference-time reasoning, remains limited to text-based problems and struggles to address tasks requiring other modalities, such as geometry.
-Overcoming these challenges calls for further advancements in the model’s cognitive capabilities.
-Appendix A Baselines
-We consider several established prompting methods as baselines as follows:
-  • Chain-of-thought (CoT) Wei et al. (2022): A prompting technique that encourages models to generate intermediate reasoning steps to enhance problem-solving capabilities.
-  • Self-Consistent Chain of Thought (SC-CoT) Wang et al.
-(2022): An extension of CoT that improves reasoning consistency by generating multiple reasoning chains and selecting the most consistent answer.
-  • Multi-Chain Reasoning (MCR) Yoran et al. (2024): enhances SC-CoT by having another LLM to assess and integrate content among the sampled reasoning chains to generate the final consistent answer.
-  • Tree of Thoughts (ToT) Yao et al.
-(2023): A method that explores multiple reasoning paths in a tree structure, allowing the model to consider various possibilities before arriving at a conclusion by tree search algorithms.
-  • Reflexion Shinn et al. (2024): A framework that enables models to reflect on their reasoning process, iteratively refining their answers based on feedback.
-  • MACM Lei et al. (2024): A multi-agent system to refine the reasoning based on iterative condition mining.
-Appendix B Computational Efficiency
-B.1 Why Measure Efficiency Through Token Usage?
-Evaluating the computational efficiency of LLMs is a critical yet nuanced task.
-A natural metric is wall-clock time, which directly reflects the real-world latency experienced during LLM inference.
-However, when interacting with LLMs primarily through API calls, as what we do in this paper, wall-clock time measurements become unreliable and difficult to interpret.
-This unreliability arises from several uncontrollable factors, including network latency, server-side load balancing, and dynamic optimizations on the provider’s infrastructure.
-These external variables introduce noise and variability that obscure the intrinsic computational characteristics of the reasoning method itself.
-To address these challenges, we adopt token usage as a more stable and conceptually meaningful proxy for computational cost.
-Token consumption directly correlates with the amount of computation performed by the LLM and is widely recognized in recent literature as a standard efficiency metric Liu et al.
-(2025); Li et al.
-(2025b).
-By focusing on token reduction, we provide a consistent basis for comparing reasoning strategies that is less susceptible to external system variability.
-B.2 Iterative Reasoning and Memory Management
-Unlike traditional CoT methods that generate reasoning process in a single pass, Meta-Reasoner pauses generation periodically to evaluate the reasoning progress and adaptively switch the reasoning strategies using a multi-armed bandit-based algorithm.
-This iterative design adds additional memory overhead due to maintaining context across multiple LLM calls.
-To mitigate this, we summarize the entire reasoning history into compact progress reports (§4.2), greatly reducing the token load per iteration.
-For example, a typical 1000-token CoT step may consume approximately 500MB of KV cache (e.g., using LLaMa-2-7B in float16 precision), while our method’s 5 iteration with 100-token steps peak at roughly 50MB each.
-Additionally, early pruning of unlikely reasoning paths further controls memory usage.
-B.3 Managing Reward Model Overhead
-To minimize the computational cost of the reward model, our design relies on summarized progress reports rather than the full reasoning history, thereby reducing token usage and associated processing overhead.
-The reward computation is further optimized by employing the LinUCB algorithm, a MAB method whose complexity scales linearly with the number of strategies.
-Since the number of strategies remains relatively small—typically between 3 and 5 in fixed settings and up to 10 to 15 in dynamic settings—the computational burden remains manageable.
-As demonstrated in Figure 2, the savings achieved by avoiding unproductive reasoning paths significantly outweigh the modest overhead introduced by the strategy selection process, resulting in an overall efficient reward computation framework.
-Appendix C Stability of Dynamic Strategy Generation
-Dynamically adding or refining strategies, as discussed in §4.3 (dynamic contextual bandit), enhances flexibility but risks instability.
-To address this trade-off, our approach incorporates several key mechanisms:
-We begin with a small set of verified strategies (Table 1) to establish a stable foundation and prevent arbitrary generation.
-Building on this, the LLM generate contextually relevant and precise strategies, like “Pause to clarify and disambiguate the current reasoning” and “Simplify by decomposing the task into smaller, manageable sub-tasks”, demonstrate how the meta-reasoner effectively targets specific reasoning bottlenecks.
-Second, we employ a contextual multi-armed bandit to balance exploration and exploitation, filtering out suboptimal strategies during LLM inference time.
-As shown in Table 6, the dynamic bandit variant achieves 89% accuracy on task Game-of-24, significantly surpassing fixed strategy sets (65%–72%).
-This highlights the dynamic bandit’s ability to prioritize effective strategies while discarding less useful ones.
-Third, the reward function described in §4.3 provides immediate feedback on strategy performance, enabling rapid deprioritization of confusing or unproductive strategies.
-Figure 3 shows a consistent increase in cumulative rewards over iterations, indicating that the system maintains coherent reasoning as it incorporates new strategies.
-Together, these mechanisms enable the dynamic contextual bandit to maintain stability without compromising adaptability.
-The observed performance gains on both the Game-of-24 and TheoremQA datasets (Table 6) show that the benefits of dynamic strategy generation substantially outweigh potential instability concerns in practice.
-Algorithm 1 Meta-Reasoner: Meta-Reasoning with Contextual Multi-Armed Bandits
-0:  LRM $M$, bandit $\mathcal{B}$, initial strategy set $\mathcal{A}_{1}$, maximum rounds $T$
-0:  Final answer $A_{\mathrm{final}}$
-1:  $C_{0}\leftarrow\emptyset$;  .$\mathcal{B}.\mathrm{Initialize}(\mathcal{A}_{1})$
-2:  $G_{0}\leftarrow$ default strategy
-3:  for $t=1$ to $T$ do
-4:     if $t>1$ then
-5:        $P_{t-1}\leftarrow f(C_{t-1})$ // Summarize the existing CoT
-6:        $x_{t-1}\leftarrow\mathrm{FeatureExtract}(P_{t-1})$ // Extract features for context
-7:        (Optional): $\mathcal{A}_{t}\leftarrow\mathcal{A}_{t-1}\cup\{\text{new strategies}\}$ // Update strategy set dynamically
-8:        $a_{t-1}\leftarrow\arg\max_{a\in\mathcal{A}_{t}}\mathrm{Score}_{\mathcal{B}}(x_{t-1},a)$ // Select strategy using bandit
-9:        $G_{t}\leftarrow a_{t-1}$ // Set current guidance
-10:     else
-11:        $G_{t}\leftarrow G_{0}$ // Use default guidance for the first iteration
-12:     end if
-13:     $s_{t}\leftarrow M(C_{t-1},G_{t})$ // Generate new CoT with integrated guidance
-14:     $C_{t}\leftarrow C_{t-1}\cup\{s_{t}\}$ // Append new reasoning step to the CoT
-15:     $r_{t}\leftarrow\mathrm{ComputeReward}(C_{t})$ // Compute reward based on the updated CoT
-16:     if $t>1$ then
-17:        .$\mathcal{B}.\mathrm{Update}(x_{t-1},a_{t-1},r_{t})$ // Update bandit with observed feedback
-18:     end if
-19:     if termination condition met then
-20:        break
-21:     end if
-22:  end for
-23:  $A_{\mathrm{final}}\leftarrow\mathrm{ExtractAnswer}(C_{t})$
-24:  return  $A_{\mathrm{final}}$
-Figure 4: Prompt Demonstration (Page-1)
-Figure 5: Prompt Demonstration (Page-2)
-Figure 6: Prompt Demonstration (Page-3)
-Figure 7: Prompt Demonstration (Page-4)